### v3

# On the universal curve with unordered marked points in positive characteristic

Abstract.
We study the relative pro-$\ell$ and continuous relative completions of the algebraic fundamental groups of universal curves over the moduli stack of curves with unordered marked points in positive characteristic. Using specialization and homotopy exact sequences, we compare the ordered and unordered settings and prove that the natural projection from the relative completion of the universal curve over the unordered moduli stack admits no section in positive characteristic. This yields a non-splitting result for the corresponding projection on algebraic fundamental groups. The present paper is a sequel to our earlier work in characteristic zero.
The first author is supported in part by National Natural Science Foundation of China (No. 12201217), by Shanghai Pilot Program for Basic Research, and by Science and Technology Commission of Shanghai Municipality (No. 22DZ2229014).
1. Introduction
Let $k$ be a field of characteristic $p>0$ and assume that 2g2+n>02g-2+n>0. Denote by ${\mathcal{M}}_{g,n/k}$ the moduli stack of smooth proper curves of genus $g$ equipped with $n$ ordered, distinct marked points over $k$, and let
${\mathcal{C}}_{g,n/k}\longrightarrow{\mathcal{M}}_{g,n/k}$
be the universal curve. The symmetric group $S_{n}$ acts freely (in the stack-theoretic sense) by permuting the marked points, which induces quotient stacks
${\mathcal{C}}_{g,[n]/k}:=[{\mathcal{C}}_{g,n/k}/S_{n}],\qquad{\mathcal{M}}_{g,[n]/k}:=[{\mathcal{M}}_{g,n/k}/S_{n}].$
We call the induced morphism
$\pi_{g,[n]/k}:{\mathcal{C}}_{g,[n]/k}\longrightarrow{\mathcal{M}}_{g,[n]/k}$
the universal curve of genus $g$ with $n$ unordered marked points over $k$.
Fix an algebraic closure $\bar{k}$ of $k$. Let $\bar{\xi}$ be a geometric point of ${\mathcal{M}}_{g,[n]/\bar{k}}$, and let $C$ denote the corresponding geometric fiber of $\pi_{g,[n]/\bar{k}}$. Choose a geometric point $\bar{y}$ of $C$. Then $\pi_{g,[n]/\bar{k}}$ induces the exact sequence of profinite groups
(1) $1\longrightarrow\pi_{1}(C,\bar{y})^{{(\ell)}}\longrightarrow\pi_{1}^{\prime}\!\bigl({\mathcal{C}}_{g,[n]/\bar{k}},\bar{y}\bigr)\longrightarrow\pi_{1}\!\bigl({\mathcal{M}}_{g,[n]/\bar{k}},\bar{\xi}\bigr)\longrightarrow 1.$
Here $\pi_{1}(C,\bar{y})^{{(\ell)}}$ denotes the pro-$\ell$ completion of $\pi_{1}(C,\bar{y})$, and $\pi_{1}^{\prime}\!\bigl({\mathcal{C}}_{g,[n]/\bar{k}},\bar{y}\bigr)$ is a canonical quotient of $\pi_{1}\!\bigl({\mathcal{C}}_{g,[n]/\bar{k}},\bar{y}\bigr)$ obtained by the SGA 1 construction (cf. [GR71, Exp. XIII, §4]). We briefly review this construction in §2.1.
Passing to continuous relative completions with respect to the $\ell$-adic monodromy representation
$\pi_{1}\!\bigl({\mathcal{C}}_{g,[n]/\bar{k}},\bar{y}\bigr)\longrightarrow\pi_{1}\!\bigl({\mathcal{M}}_{g,[n]/\bar{k}},\bar{\xi}\bigr)\longrightarrow\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell}}),$
where $\ell\neq\operatorname{char}(k)$ and $H_{{\mathbb{Q}}_{\ell}}:=H^{1}_{\mathrm{\acute{e}t}}(C,{{\mathbb{Q}}_{\ell}}(1))$, one obtains an exact sequence of pro-algebraic $\mathbb{Q}$$_{\ell}$–groups (see §5)
(2) $1\longrightarrow F^{\mathrm{un}}_{g}\longrightarrow{\mathcal{G}}_{{\mathcal{C}}_{g,[n]},\bar{\xi}}\longrightarrow{\mathcal{G}}_{g,[n],\bar{\xi}}\longrightarrow 1.$
Here $F^{\mathrm{un}}_{g}$ is the $\ell$-adic unipotent completion of $\pi_{1}(C,\bar{y})$.
Theorem 1.
Suppose that $\operatorname{char}(k)=p>0$, $\ell\not=p$ an odd prime, and $g\geq 3$. Then the exact sequence (2) does not admit any section.
Corollary 2.
If $g\geq 3$, the canonical surjection
$\pi_{1}\!\bigl({\mathcal{C}}_{g,[n]/k},\bar{y}\bigr)\longrightarrow\pi_{1}\!\bigl({\mathcal{M}}_{g,[n]/k},\bar{\xi}\bigr)$
admits no continuous section. In particular, the universal curve
${\mathcal{C}}_{g,[n]/k}\longrightarrow{\mathcal{M}}_{g,[n]/k}$
has no section.
This paper is a sequel to [LW25], where the corresponding nonsplitting results for universal curves with unordered marked points were established in characteristic 0. Our aim here is to extend those results to positive characteristic. Although the overall strategy follows the same general framework as in the characteristic 0 case, carrying out this extension requires substantial additional work. In characteristic 0, the essential geometric information needed for the argument is encoded in the weight filtration on relative and continuous relative completions. In positive characteristic, while continuous relative completion remains available for algebraic fundamental groups, transferring the corresponding structural properties across characteristics is far from automatic. To address this, we make systematic use of relative pro-$\ell$ completion and continuous relative completion, together with comparison isomorphisms between characteristic 0 and characteristic $p$ established in [WAT19], which allow us to transport the necessary monodromy and homotopy-theoretic information into characteristic $p$. These tools play a central role throughout the paper and are indispensable for the arguments that follow. Familiarity with the methods developed in [LW25] will therefore be beneficial to the reader.
2. Moduli of Curves with unordered marked points
Let $k$ be a field of characteristic $p>0$, and fix an algebraic closure $\bar{k}\subset\Omega$, where $\Omega$ is a chosen algebraically closed field containing $k$. Assume throughout that $g,n\geq 0$ and 2g2+n>02g-2+n>0.
Over $\operatorname{Spec}{\mathbb{Z}}$, let ${\mathcal{M}}_{g,n/{\mathbb{Z}}}$ denote the moduli stack of smooth, proper curves of genus $g$ with $n$ ordered distinct marked points. Its base change
${\mathcal{M}}_{g,n/k}:={\mathcal{M}}_{g,n/{\mathbb{Z}}}\times_{\operatorname{Spec}{\mathbb{Z}}}\operatorname{Spec}k$
is a smooth Deligne–Mumford stack over $k$.
The symmetric group $S_{n}$ defines a constant finite étale group scheme over $\operatorname{Spec}{\mathbb{Z}}$. By base change,
$S_{n/k}:=S_{n}\times_{\operatorname{Spec}{\mathbb{Z}}}\operatorname{Spec}k$
is a constant finite étale group scheme over $k$ acting on ${\mathcal{M}}_{g,n/k}$ by permuting the marked points. Since ${\mathcal{M}}_{g,n/k}$ is a smooth Deligne–Mumford stack, the quotient stack
${\mathcal{M}}_{g,[n]/k}:=[\,{\mathcal{M}}_{g,n/k}/S_{n/k}\,]$
is also a smooth Deligne–Mumford stack over $k$. It parametrizes smooth, proper genus $g$ curves with $n$ unordered marked points.
The universal curve over ${\mathcal{M}}_{g,n/{\mathbb{Z}}}$ base changes to the universal curve over $k$:
$\pi_{g,n/k}:{\mathcal{C}}_{g,n/k}\longrightarrow{\mathcal{M}}_{g,n/k}.$
The $S_{n/k}$-action on ${\mathcal{M}}_{g,n/k}$ lifts to ${\mathcal{C}}_{g,n/k}$, and the resulting quotient stack
${\mathcal{C}}_{g,[n]/k}:=[\,{\mathcal{C}}_{g,n/k}/S_{n/k}\,]$
is a Deligne–Mumford stack with induced morphism
$\pi_{g,[n]/k}:{\mathcal{C}}_{g,[n]/k}\longrightarrow{\mathcal{M}}_{g,[n]/k},$
which we refer to as the universal curve of genus $g$ with $n$ unordered marked points over $k$. Both projections
${\mathcal{M}}_{g,n/k}\longrightarrow{\mathcal{M}}_{g,[n]/k}\qquad\text{and}\qquad{\mathcal{C}}_{g,n/k}\longrightarrow{\mathcal{C}}_{g,[n]/k}$
are finite étale $S_{n/k}$-torsors.
Finally, for any connected Deligne–Mumford stack $\mathcal{X}$} over a field $L$, we write $\pi_{1}({\mathcal{X}}_{/L})$ for its (profinite) algebraic fundamental group in the sense of SGA 1. When working over $\mathbb{C}$}, we instead use the notation $\pi_{1}^{\mathrm{top}}$ and $\pi_{1}^{\mathrm{orb}}$ for the topological and orbifold fundamental groups, respectively.
2.1. Homotopy exact sequences in positive characteristic
In this subsection, we briefly recall the SGA 1 construction of the homotopy exact sequence for a proper smooth morphism with geometrically connected fibers, and then apply it to the ordered and unordered universal curves (cf. [GR71, SGA 1, Exp. XIII, §4]; see also [WAT19, Prop. 4.2]).
Let $L$ be a field, and let $T$ be a connected, locally noetherian Deligne–Mumford stack over $L$. Let $f:C\longrightarrow T$ be a proper smooth morphism whose geometric fibers are connected curves of genus $g\geq 2$. Fix geometric points $\bar{\zeta}$ in $T$ and $\bar{a}$ in the fiber $C_{\bar{\zeta}}$ of $f$ over $\bar{\zeta}$. Let $\ell$ be a prime number different from $\operatorname{char}(k)$. Following [GR71, SGA 1, Exp. XIII, §4.3–4.4], set
$K:=\ker\!\bigl(\pi_{1}(C,\bar{a})\longrightarrow\pi_{1}(T,\bar{\zeta})\bigr),\qquad K^{(\ell)}:=\text{maximal pro-}\ell\text{ quotient of }K,$
and let
$N:=\ker\!\bigl(K\longrightarrow K^{(\ell)}\bigr).$
Then $N$ is a distinguished closed normal subgroup of $\pi_{1}(C,\bar{a})$, and we define
$\pi_{1}^{\prime}(C,\bar{a}):=\pi_{1}(C,\bar{a})/N.$
We also write $\pi_{1}(C_{\bar{\zeta}},\bar{a})^{(\ell)}$ for the maximal pro-$\ell$ quotient of $\pi_{1}(C_{\bar{\zeta}},\bar{a})$. With this notation, [WAT19, Prop. 4.2] yields the exact sequence
(3) $1\longrightarrow\pi_{1}(C_{\bar{\zeta}},\bar{a})^{(\ell)}\longrightarrow\pi_{1}^{\prime}(C,\bar{a})\longrightarrow\pi_{1}(T,\bar{\zeta})\longrightarrow 1.$
Assume $g\geq 2$, and let $L$ be a field of characteristic $p\geq 0$. Consider the ordered and unordered universal curves
$\pi_{g,n/L}:{\mathcal{C}}_{g,n/L}\longrightarrow{\mathcal{M}}_{g,n/L},\qquad\pi_{g,[n]/L}:{\mathcal{C}}_{g,[n]/L}\longrightarrow{\mathcal{M}}_{g,[n]/L}.$
Fix a geometric point $\bar{\zeta}$ of ${\mathcal{M}}_{g,n/L}$, which we also view as a geometric point of ${\mathcal{M}}_{g,[n]/L}$ via the finite étale projection ${\mathcal{M}}_{g,n/L}\to{\mathcal{M}}_{g,[n]/L}$. Let $C_{\bar{\zeta}}$ denote the corresponding geometric fiber, and choose a geometric point $\bar{a}\in C_{\bar{\zeta}}$. Via the natural morphisms
$C_{\bar{\zeta}}\longrightarrow{\mathcal{C}}_{g,n/L}\longrightarrow{\mathcal{C}}_{g,[n]/L},$
we regard $\bar{a}$ as a geometric point of both ${\mathcal{C}}_{g,n/L}$ and ${\mathcal{C}}_{g,[n]/L}$.
Applying the SGA 1 construction (3) to $\pi_{g,n/L}$ yields an exact sequence
(4) $1\longrightarrow\pi_{1}(C_{\bar{\zeta}},\bar{a})^{(\ell)}\longrightarrow\pi_{1}^{\prime}\bigl({\mathcal{C}}_{g,n/L},\bar{a}\bigr)\longrightarrow\pi_{1}\bigl({\mathcal{M}}_{g,n/L},\bar{\zeta}\bigr)\longrightarrow 1.$
Likewise, applying the same construction to the unordered universal curve $\pi_{g,[n]/L}$ gives an exact sequence
(5) $1\longrightarrow\pi_{1}(C_{\bar{\zeta}},\bar{a})^{(\ell)}\longrightarrow\pi_{1}^{\prime}\bigl({\mathcal{C}}_{g,[n]/L},\bar{a}\bigr)\longrightarrow\pi_{1}\bigl({\mathcal{M}}_{g,[n]/L},\bar{\zeta}\bigr)\longrightarrow 1.$
Passing from ordered to unordered moduli is induced by the finite étale $S_{n}$-torsors ${\mathcal{M}}_{g,n/L}\to{\mathcal{M}}_{g,[n]/L}$ and ${\mathcal{C}}_{g,n/L}\to{\mathcal{C}}_{g,[n]/L}$. These morphisms induce maps between the exact sequences (4) and (5), summarized in the following commutative diagram:
[…]
The vertical arrows are induced by the finite étale quotient maps associated with the $S_{n}$-action. In particular, the right-hand square of Diagram (6) is a pullback square in the category of profinite groups.
3. Configuration spaces and their fundamental groups
Let $L$ be an algebraically closed field and let $C$ be a proper smooth connected curve over $L$. For a positive integer $n$, the ordered configuration space of $n$ points on $C$ is the open subscheme
$F(C,n):=C^{n}\setminus\Delta,$
where $\Delta\subset C^{n}$ denotes the fat diagonal
$\Delta:=\{(x_{1},\dots,x_{n})\in C^{n}\mid x_{i}=x_{j}\text{ for some }i\neq j\}.$
Thus $F(C,n)$ parametrizes $n$-tuples of pairwise distinct points of $C$.
The symmetric group $S_{n}$ acts freely on $F(C,n)$ by permuting the ordered points, and the quotient
$F(C,[n]):=F(C,n)/S_{n}$
is the unordered configuration space, parametrizing unordered $n$-tuples of distinct points of $C$. The projection
$F(C,n)\longrightarrow F(C,[n])$
is therefore a finite étale $S_{n}$-torsor.
In particular, for a geometric point $\bar{a}$ of $F(C,n)$ (also viewed as a geometric point in $F(C,[n])$), one obtains an exact sequence of étale fundamental groups
$1\longrightarrow\pi_{1}\bigl(F(C,n),\bar{a}\bigr)\longrightarrow\pi_{1}\bigl(F(C,[n]),\bar{a}\bigr)\longrightarrow S_{n}\longrightarrow 1.$
3.1. The pro-$\ell$ completion of $\pi_{1}(F(C,n))$
Suppose that $k$ is a field of characteristic $p>0$, and let $\bar{k}$ be an algebraic closure of $k$. Let $C$ be a connected, proper, smooth curve of genus $g\geq 3$ over $\bar{k}$. Choose a complete DVR $R$ of characteristic 0 with residue field $\bar{k}$ and a proper, smooth morphism
$X\longrightarrow\operatorname{Spec}R$
whose special fiber is isomorphic to $C$. (For example, one may take $R=W(\bar{k})$, the ring of Witt vectors, which is a complete DVR of mixed characteristic $(0,p)$.)
Let $K$ be the fraction field of $R$ and $\bar{K}$ an algebraic closure of $K$. For an integer $n\geq 1$, let $F(X,n)$ denote the ordered configuration space of $X$ with $n$ labeled points. The Fulton–MacPherson compactification [FM94] of $F(X,n)$ is proper and smooth over $R$, and $F(X,n)$ is the complement of a relative normal crossings divisor.
Let $\bar{\eta}:\operatorname{Spec}\bar{K}\to\operatorname{Spec}R$ and $\bar{\xi}:\operatorname{Spec}\bar{k}\to\operatorname{Spec}R$ be the geometric points corresponding to $\bar{K}$ and $\bar{k}$, respectively, and note that $X_{\bar{\xi}}\cong C$. Fix an embedding $\iota:\bar{K}\hookrightarrow{\mathbb{C}}$, and denote the associated complex curve of $X_{\bar{\eta}}$ by
$C^{\mathrm{an}}:=X_{\bar{\eta}}\otimes_{\bar{K},\iota}{\mathbb{C}}.$
Whenever we use an analytic description, we implicitly pass through such a choice of $\iota$. Then
$F(X_{\bar{\eta}},n)\;\cong\;F(X,n)_{\bar{K}}\quad\text{and}\quad F(C,n)\;\cong\;F(X,n)_{\bar{k}}.$
Let $\bar{x}$ and $\bar{y}$ be geometric points of $F(X_{\bar{\eta}},n)$ and $F(C,n)$, respectively. Via the natural morphisms $F(X_{\bar{\eta}},n)\longrightarrow F(X,n)$ and $F(C,n)\longrightarrow F(X,n)$, we regard $\bar{x}$ and $\bar{y}$ as geometric points of $F(X,n)$. Fixing an isomorphism
$\lambda:\pi_{1}\bigl(F(X,n),\bar{x}\bigr)\xrightarrow{\sim}\pi_{1}\bigl(F(X,n),\bar{y}\bigr),$
we obtain a specialization homomorphism
$\mathrm{sp}_{\ell}:\pi_{1}\bigl(F(X_{\bar{\eta}},n),\bar{x}\bigr)^{{(\ell)}}\;\longrightarrow\;\pi_{1}\bigl(F(C,n),\bar{y}\bigr)^{{(\ell)}},$
for each prime $\ell\neq p$. By [GR71, Exposé XIII, Thm. 2.10, Cor. 2.9] and the fact that the Fulton–MacPherson compactification is smooth over $R$ and that $F(X,n)$ is the complement of a relative normal crossings divisor, this specialization map $\mathrm{sp}_{\ell}$ is an isomorphism.
Denote $\pi_{1}^{\mathrm{top}}(F(C^{\mathrm{an}},n))$ by $P_{g,n}$. Then $\pi_{1}\bigl(F(X_{\bar{\eta}},n),\bar{x}\bigr)$ is naturally isomorphic to the profinite completion $\widehat{P}_{g,n}$, and hence we obtain an isomorphism
(7) $\pi_{1}\bigl(F(C,n),\bar{y}\bigr)^{{(\ell)}}\;\cong\;P_{g,n}^{{(\ell)}}.$
3.2. The $\ell$-adic unipotent completion
Let $\ell\neq p$ be a prime. Fix a geometric point $\bar{y}$ of $F(C,n)$ (also view it as a geometric point of $F(C,[n])$) and for simplicity write
$F_{g,n}:=\pi_{1}\bigl(F(C,n),\bar{y}\bigr),\qquad F_{g,[n]}:=\pi_{1}\bigl(F(C,[n]),\bar{y}\bigr).$
Note $F_{g,1}=\pi_{1}(C,\bar{y})$. Denote by $F_{g,n}^{{(\ell)}}$ and $F_{g,[n]}^{{(\ell)}}$ their pro-$\ell$ completions:
$F_{g,n}^{{(\ell)}}:=\pi_{1}\bigl(F(C,n),\bar{y}\bigr)^{(\ell)},\qquad F_{g,[n]}^{{(\ell)}}:=\pi_{1}\bigl(F(C,[n]),\bar{y}\bigr)^{(\ell)}.$
We now pass to the $\ell$-adic unipotent completions. Following [HM03, Appendix A], let $F_{g,n}^{\mathrm{un}}$ and $F_{g,[n]}^{\mathrm{un}}$ denote the $\ell$-adic unipotent completions of $F_{g,n}$ and $F_{g,[n]}$ over $\mathbb{Q}$$_{\ell}$, respectively. These are pro-unipotent algebraic $\mathbb{Q}$$_{\ell}$}-groups equipped with Zariski-dense continuous homomorphisms
$F_{g,n}\longrightarrow F_{g,n}^{\mathrm{un}}({\mathbb{Q}}_{\ell}),\qquad F_{g,[n]}\longrightarrow F_{g,[n]}^{\mathrm{un}}({\mathbb{Q}}_{\ell}).$
[…]
$F_{g,n}^{{(\ell)}}\longrightarrow F_{g,n}^{\mathrm{un}}({\mathbb{Q}}_{\ell}),\qquad F_{g,[n]}^{{(\ell)}}\longrightarrow F_{g,[n]}^{\mathrm{un}}({\mathbb{Q}}_{\ell})$
are again Zariski-dense. We define their Lie algebras
${\mathfrak{p}}_{g,n}:=\operatorname{Lie}\bigl(F_{g,n}^{\mathrm{un}}\bigr),\qquad{\mathfrak{p}}_{g,[n]}:=\operatorname{Lie}\bigl(F_{g,[n]}^{\mathrm{un}}\bigr),$
which are pronilpotent Lie algebras over $\mathbb{Q}$$_{\ell}$. When $n=1$, we write $F_{g}^{\mathrm{un}}$ and ${\mathfrak{p}}_{g}$ for $F_{g,1}^{\mathrm{un}}$ and ${\mathfrak{p}}_{g,1}$.
For a group $\Gamma$} and a field $L$, set $H_{1}({\Gamma},L):={\Gamma}^{\mathrm{ab}}\otimes_{\mathbb{Z}}L$.
Lemma 3.1.
[…]
$H^{1}_{\mathrm{\acute{e}t}}(C,{\mathbb{Q}}_{\ell}(1))\;\cong\;H_{1}\bigl(\pi_{1}(C)^{{(\ell)}},{\mathbb{Q}}_{\ell}\bigr),$
compatible with the natural symplectic form on the left-hand side.
Proof.
Since $C$ is proper and smooth over $L$, the comparison between étale cohomology of $C$ and continuous cohomology of its profinite fundamental group yields
$H^{1}_{\mathrm{\acute{e}t}}(C,{\mathbb{Q}}_{\ell}(1))\;\cong\;H^{1}_{\mathrm{cts}}\bigl(\pi_{1}(C),{\mathbb{Q}}_{\ell}(1)\bigr)\;\cong\;\operatorname{Hom}_{\mathrm{cts}}\!\bigl(\pi_{1}(C),{\mathbb{Q}}_{\ell}(1)\bigr).$
Since any continuous map $\pi_{1}(C)\to{\mathbb{Q}}_{\ell}(1)$ factors through the pro-$\ell$ abelianization of $\pi_{1}(C)$, we obtain canonical isomorphisms
$\operatorname{Hom}_{\mathrm{cts}}\!\bigl(\pi_{1}(C),{\mathbb{Q}}_{\ell}(1)\bigr)$ $\cong\operatorname{Hom}_{\mathrm{cts}}\!\bigl(\pi_{1}(C)^{{(\ell)}},{\mathbb{Q}}_{\ell}(1)\bigr)$
$\cong\operatorname{Hom}_{{\mathbb{Q}}_{\ell}}\!\bigl(H_{1}(\pi_{1}(C)^{{(\ell)}},{\mathbb{Q}}_{\ell}),{\mathbb{Q}}_{\ell}(1)\bigr).$
Since $H_{1}(\pi_{1}(C)^{{(\ell)}},{\mathbb{Q}}_{\ell})\cong{\mathbb{Q}}_{\ell}^{2g}$ is finite-dimensional, the last space is canonically isomorphic to $H_{1}(\pi_{1}(C)^{{(\ell)}},{\mathbb{Q}}_{\ell})^{\vee}\otimes{\mathbb{Q}}_{\ell}(1).$ On the other hand, there is a perfect alternating pairing
$H^{1}_{\mathrm{\acute{e}t}}(C,{\mathbb{Q}}_{\ell}(1))\otimes H^{1}_{\mathrm{\acute{e}t}}(C,{\mathbb{Q}}_{\ell}(1))\longrightarrow H^{2}_{\mathrm{\acute{e}t}}(C,{\mathbb{Q}}_{\ell}(2))\cong{\mathbb{Q}}_{\ell}(1),$
which induces a canonical isomorphism $H^{1}_{\mathrm{\acute{e}t}}(C,{\mathbb{Q}}_{\ell}(1))\xrightarrow{\sim}H^{1}_{\mathrm{\acute{e}t}}(C,{\mathbb{Q}}_{\ell}(1))^{\vee}(1).$ Combining this with the identification above, we compute
$H^{1}_{\mathrm{\acute{e}t}}(C,{\mathbb{Q}}_{\ell}(1))$ $\cong\bigl(H_{1}(\pi_{1}(C)^{{(\ell)}},{\mathbb{Q}}_{\ell})^{\vee}\otimes{\mathbb{Q}}_{\ell}(1)\bigr)^{\vee}\otimes{\mathbb{Q}}_{\ell}(1)$
$\cong H_{1}(\pi_{1}(C)^{{(\ell)}},{\mathbb{Q}}_{\ell})\otimes{\mathbb{Q}}_{\ell}(-1)\otimes{\mathbb{Q}}_{\ell}(1)$
$\cong H_{1}(\pi_{1}(C)^{{(\ell)}},{\mathbb{Q}}_{\ell}).$
3.3. Abelianizations of unipotent completions of configuration spaces
We study the abelianizations $H_{1}({\mathfrak{p}}_{g,n})$ and $H_{1}({\mathfrak{p}}_{g,[n]})$ via the comparison with the characteristic zero case. Set $H_{{\mathbb{Q}}_{\ell}}:=H^{1}_{\mathrm{\acute{e}t}}(C,{\mathbb{Q}}_{\ell}(1))$. The $S_{n}$-action on $F(C,n)$ by permuting points induces an action on homology. The following proposition confirms that the familiar characteristic-zero behavior holds equally in positive characteristic.
Proposition 3.2.
For $g\geq 0$ and $n\geq 1$, there are natural isomorphisms
$H_{1}({\mathfrak{p}}_{g,n})\;\cong\;H_{1}\bigl(F_{g,n}^{{(\ell)}},{\mathbb{Q}}_{\ell}\bigr)\;\cong\;H_{{\mathbb{Q}}_{\ell}}^{\oplus n}.$
Proof.
[…]
$H_{1}\bigl(P_{g,n},{\mathbb{Z}}\bigr)\;\cong\;H_{1}\bigl(F(C^{\mathrm{an}},n),{\mathbb{Z}}\bigr)\;\cong\;H_{1}(C^{\mathrm{an}},{\mathbb{Z}})^{\oplus n}.$
Since $P_{g,n}$ is finitely generated, tensoring with $\mathbb{Q}$$_{\ell}$, we obtain
$H_{1}\bigl(P_{g,n},{\mathbb{Q}}_{\ell}\bigr)\;\cong\;H_{1}\bigl(P_{g,n}^{(\ell)},{\mathbb{Q}}_{\ell}\bigr)\;\cong\;H_{1}(C^{\mathrm{an}},{\mathbb{Q}}_{\ell})^{\oplus n}.$
Finally, since $\pi_{1}^{\mathrm{top}}(C^{\mathrm{an}})\textcent$ is finitely generated, combining the specialization isomorphism $\pi_{1}(C)^{{(\ell)}}\cong\pi_{1}^{\mathrm{top}}(C^{\mathrm{an}})^{(\ell)}$ with Lemma 3.1, we have natural isomorphisms
$H_{1}(C^{\mathrm{an}},{\mathbb{Q}}_{\ell})\;\cong\;H_{1}\bigl(\pi_{1}^{\mathrm{top}}(C^{\mathrm{an}})^{{(\ell)}},{\mathbb{Q}}_{\ell}\bigr)\;\cong\;H_{1}\bigl(\pi_{1}(C)^{{(\ell)}},{\mathbb{Q}}_{\ell}\bigr)\;\cong\;H_{{\mathbb{Q}}_{\ell}},$
and therefore
$H_{1}\bigl(F_{g,n}^{{(\ell)}},{\mathbb{Q}}_{\ell}\bigr)\cong H_{1}\bigl(P_{g,n}^{{(\ell)}},{\mathbb{Q}}_{\ell}\bigr)\;\cong\;H_{{\mathbb{Q}}_{\ell}}^{\oplus n}.$
[…]
$H_{1}({\mathfrak{p}}_{g,n})\cong H_{1}\bigl(\operatorname{Lie}(F^{{(\ell)}\mathrm{un}}_{g,n})\bigr)\cong H_{1}\bigl(F^{(\ell)}_{g,n},{{\mathbb{Q}}_{\ell}}\bigr)\cong H_{{\mathbb{Q}}_{\ell}}^{\oplus n}.$
The previous result computes the first homology in the ordered case. We now pass to the unordered configuration space by taking $S_{n}$-coinvariants.
Proposition 3.3.
For $g\geq 0$ and $n\geq 1$, there are natural isomorphisms
$H_{1}({\mathfrak{p}}_{g,[n]})\;\cong\;H_{1}\bigl(F_{g,[n]}^{{(\ell)}},{\mathbb{Q}}_{\ell}\bigr)\;\cong\;\bigl(H_{{\mathbb{Q}}_{\ell}}^{\oplus n}\bigr)_{S_{n}}\;\cong\;H_{{\mathbb{Q}}_{\ell}}.$
Proof.
[…]
$H_{1}({\mathfrak{p}}_{g,[n]})\;\cong\;H_{1}\bigl(F_{g,[n]}^{{(\ell)}},{\mathbb{Q}}_{\ell}\bigr).$
The finite étale Galois cover
$F(C,n)\longrightarrow F(C,[n])$
with group $S_{n}$ induces an exact sequence of algebraic fundamental groups
$1\longrightarrow F_{g,n}\longrightarrow F_{g,[n]}\longrightarrow S_{n}\longrightarrow 1.$
The Hochschild–Serre spectral sequence for group homology with coefficients in $\mathbb{Q}$$_{\ell}$ has
$E^{2}_{p,q}=H_{p}\bigl(S_{n},\,H_{q}(F_{g,n},{\mathbb{Q}}_{\ell})\bigr)\;\Longrightarrow\;H_{p+q}\bigl(F_{g,[n]},{\mathbb{Q}}_{\ell}\bigr).$
Since $S_{n}$ is finite and $|S_{n}|$ is invertible in $\mathbb{Q}$$_{\ell}$, one has $H_{p}(S_{n},V)=0$ for all $p>0$ and all ${\mathbb{Q}}_{\ell}[S_{n}]$-modules $V$. In degree 1 this yields a canonical isomorphism
$H_{1}\bigl(F_{g,[n]},{\mathbb{Q}}_{\ell}\bigr)\;\cong\;H_{0}\bigl(S_{n},\,H_{1}(F_{g,n},{\mathbb{Q}}_{\ell})\bigr)=H_{1}(F_{g,n},{\mathbb{Q}}_{\ell})_{S_{n}}.$
Indeed, for a profinite group $G$, the abelianization commutes with passage to the maximal pro-$\ell$ quotient, and tensoring with $\mathbb{Q}$$_{\ell}$ kills all prime-to-$\ell$ components. Consequently,
$G^{\mathrm{ab}}\otimes{\mathbb{Q}}_{\ell}\;\cong\;(G^{\mathrm{ab}})^{{(\ell)}}\otimes{\mathbb{Q}}_{\ell}\;\cong\;(G^{{(\ell)}})^{\mathrm{ab}}\otimes{\mathbb{Q}}_{\ell},$
and the natural homomorphism $G\to G^{{(\ell)}}$ induces an isomorphism on $H_{1}(-,{\mathbb{Q}}_{\ell})$. Thus
$H_{1}\bigl(F_{g,[n]}^{{(\ell)}},{\mathbb{Q}}_{\ell}\bigr)\;\cong\;H_{1}\bigl(F_{g,n}^{{(\ell)}},{\mathbb{Q}}_{\ell}\bigr)_{S_{n}}.$
By the computation in the ordered case (Proposition 3.2),
$H_{1}\bigl(F_{g,n}^{{(\ell)}},{\mathbb{Q}}_{\ell}\bigr)\;\cong\;H_{{\mathbb{Q}}_{\ell}}^{\oplus n},$
where $S_{n}$ acts by permuting the $n$ summands. Hence
$H_{1}\bigl(F_{g,[n]}^{{(\ell)}},{\mathbb{Q}}_{\ell}\bigr)\;\cong\;\bigl(H_{{\mathbb{Q}}_{\ell}}^{\oplus n}\bigr)_{S_{n}}.$
Finally, the $S_{n}$-coinvariants of $H_{{\mathbb{Q}}_{\ell}}^{\oplus n}$ are canonically isomorphic to $H_{{\mathbb{Q}}_{\ell}}$. Combining the above identifications gives the natural isomorphisms
$H_{1}({\mathfrak{p}}_{g,[n]})\;\cong\;H_{1}\bigl(F_{g,[n]}^{{(\ell)}},{\mathbb{Q}}_{\ell}\bigr)\;\cong\;\bigl(H_{{\mathbb{Q}}_{\ell}}^{\oplus n}\bigr)_{S_{n}}\;\cong\;H_{{\mathbb{Q}}_{\ell}},$
which proves the proposition. ∎
4. Relative Pro-$\ell$ Completion
A key ingredient in this work is the study of the algebraic fundamental group $\pi_{1}({\mathcal{M}}_{g/\bar{k}})$ for a field $k$ of characteristic $p>0$. A natural first idea is to compare it with the mapping class group
${\Gamma}_{g}\;\cong\;\pi_{1}^{\mathrm{orb}}({\mathcal{M}}_{g/{\mathbb{C}}})$
via pro-$\ell$ completion. However, when $g\geq 3$, the mapping class group ${\Gamma}_{g}$ is perfect, and hence its pro-$\ell$ completion is trivial. As a result, the usual pro-$\ell$ completion loses all information and cannot be used to transfer the structure of $\pi_{1}({\mathcal{M}}_{g/{\mathbb{C}}})$ to the study of $\pi_{1}({\mathcal{M}}_{g/\bar{k}})$.
To overcome this, one instead considers the relative pro-$\ell$ completion. This construction retains essential information about the mapping class group by taking into account the natural $\ell$-adic monodromy representation $\rho:{\Gamma}_{g}\longrightarrow\operatorname{Sp}_{2g}({\mathbb{Z}}_{\ell})$ and completing ${\Gamma}_{g}$ in a manner that is compatible with $\rho$. In this way, the resulting relative pro-$\ell$ completion preserves the $\ell$-adic structure coming from the geometry of ${\mathcal{M}}_{g}$. As shown in [HM09], the resulting relative pro-$\ell$ completion is nontrivial and captures the geometric structure of ${\mathcal{M}}_{g}$ much more effectively than the ordinary pro-$\ell$ completion. For instance, when marked points are present, the mapping class group embeds into its relative pro-$\ell$ completion ([HM09, Prop. 3.1]).
Because the second author applied this approach in the positive characteristic setting in [WAT19], we briefly recall below the definition and the basic properties of relative pro-$\ell$ completion. Full details and extensive background can be found in [HM09].
Definition 4.1 (Relative pro-$\ell$ completion).
Let $\Gamma$ be a discrete group or a profinite group and let
$\rho:\Gamma\longrightarrow P$
be a continuous homomorphism into a profinite group $P$ whose image is dense. The relative pro-$\ell$ completion of $\Gamma$ with respect to $\rho$ is a profinite group $\Gamma^{{\mathrm{rel}(\ell)}}$ equipped with a continuous homomorphism $\rho^{(\ell)}:\Gamma\to\Gamma^{{\mathrm{rel}(\ell)}}$ satisfying the following properties:
  1. (1)
    there is an exact sequence of profinite groups
    $1\longrightarrow K\longrightarrow\Gamma^{{\mathrm{rel}(\ell)}}\xrightarrow{\;\tilde{\rho}\;}P\longrightarrow 1,$
    where $K$ is a pro-$\ell$ group and the map $\tilde{\rho}:\Gamma^{{\mathrm{rel}(\ell)}}\to P$ is induced by $\rho$ such that $\tilde{\rho}\circ\rho^{(\ell)}=\rho$;
  2. (2)
    the image $\rho^{(\ell)}(\Gamma)$ is dense in $\Gamma^{{\mathrm{rel}(\ell)}}$;
  3. (3)
    $\Gamma^{{\mathrm{rel}(\ell)}}$ is universal with these properties: if $G$ is a profinite group equipped with a surjection $G\to P$ whose kernel is pro-$\ell$ and if $\varphi:\Gamma\to G$ is a continuous homomorphism lifting $\rho$, then there exists a unique continuous homomorphism $\tilde{\varphi}:\Gamma^{{\mathrm{rel}(\ell)}}\to G$ such that $\varphi=\tilde{\varphi}\circ\rho^{(\ell)}$.
4.1. The relative pro-$\ell$ completion of the mapping class group
Let $L$ be a field of characteristic $p\geq 0$, and fix an algebraic closure $\bar{L}$ of $L$. Let $\pi:=\pi_{g,n/\bar{L}}:{\mathcal{C}}_{g,n/\bar{L}}\longrightarrow{\mathcal{M}}_{g,n/\bar{L}}$ denote the universal curve. Fix a prime $\ell\neq p$, and consider the $\ell$-adic local system
$\mathbb{H}_{{\mathbb{Z}}_{\ell}}:=R^{1}\pi_{*}{\mathbb{Z}}_{\ell}(1)$
on ${\mathcal{M}}_{g,n/\bar{L}}$. For a geometric point $\bar{\zeta}$ of ${\mathcal{M}}_{g,n/\bar{L}}$, let $C_{\bar{\zeta}}$ be the corresponding geometric fiber of $\pi$, and set $H_{{\mathbb{Z}}_{\ell}}:=\mathbb{H}_{{\mathbb{Z}}_{\ell},\bar{\zeta}}=H^{1}_{\mathrm{\acute{e}t}}(C_{\bar{\zeta}},{\mathbb{Z}}_{\ell}(1))$. The natural cup product pairing endows $H_{{\mathbb{Z}}_{\ell}}$ with a symplectic structure, yielding a monodromy representation
$\rho_{{\mathbb{Z}}_{\ell},\bar{\zeta}}:\pi_{1}({\mathcal{M}}_{g,n/\bar{L}},\bar{\zeta})\longrightarrow\operatorname{Sp}(H_{{\mathbb{Z}}_{\ell}}).$
When $\operatorname{char}(L)=p>0$ and $\ell\neq p$, the second author proved in [WAT19, Thm. 5.6] that the relative pro-$\ell$ completion of $\pi_{1}({\mathcal{M}}_{g,n/\bar{L}},\bar{\zeta})$ with respect to $\rho_{{\mathbb{Z}}_{\ell},\bar{\zeta}}$ is naturally isomorphic to ${\Gamma}_{g,n}^{{\mathrm{rel}(\ell)}}$, the relative pro-$\ell$ completion of the mapping class group ${\Gamma}_{g,n}$ with respect to its standard symplectic representation
${\Gamma}_{g,n}\longrightarrow\operatorname{Sp}\bigl(H_{1}(\Sigma_{g},{\mathbb{Z}}_{\ell})\bigr),$
where $\Sigma_{g}$ is a compact oriented surface of genus $g$.
4.2. Specialization for the universal curve in the ordered case
Let $k$ be a field of characteristic $p>0$ with algebraic closure $\bar{k}$. Let $R:=W(\bar{k})$ be the ring of Witt vectors of $\bar{k}$, let $K:=\mathrm{Frac}(R)$ be its fraction field, and fix an algebraic closure $\bar{K}$ of $K$. Let $\bar{\xi}$ and $\bar{\eta}$ be geometric points of $\mathcal{M}_{g,n/\bar{k}}$ and $\mathcal{M}_{g,n/\bar{K}}$, respectively. Via the canonical morphisms $\mathcal{M}_{g,n/\bar{k}}\to\mathcal{M}_{g,n/R}$ and $\mathcal{M}_{g,n/\bar{K}}\to\mathcal{M}_{g,n/R}$, we also regard $\bar{\xi}$ and $\bar{\eta}$ as geometric points of $\mathcal{M}_{g,n/R}$. Let $\pi_{g,n/R}:\mathcal{C}_{g,n/R}\to\mathcal{M}_{g,n/R}$ be the universal curve, and denote by $C_{\bar{\xi}}$ and $C_{\bar{\eta}}$ its fibers over $\bar{\xi}$ and $\bar{\eta}$, respectively. Fix geometric points $\bar{y}\in C_{\bar{\xi}}$ and $\bar{x}\in C_{\bar{\eta}}$.
Proposition 4.2.
For $g\geq 2$ and $n\geq 0$, there is a commutative diagram:
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(C_{\bar{\xi}},\bar{y})^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{C}_{g,n/\bar{k}},\bar{y})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g,n/\bar{k}},\bar{\xi})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(C_{\bar{\eta}},\bar{x})^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\pi_{1}(\mathcal{C}_{g,n/\bar{K}},\bar{x})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\pi_{1}(\mathcal{M}_{g,n/\bar{K}},\bar{\eta})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${1,}$
where the rows are exact and all vertical maps are isomorphisms.
Proof.
Fix an isomorphism
$\lambda:\pi_{1}(\mathcal{C}_{g,n/R},\bar{x})\;\xrightarrow{\;\sim\;}\pi_{1}(\mathcal{C}_{g,n/R},\bar{y}).$
This induces isomorphisms $\lambda^{\prime}:\pi_{1}^{\prime}(\mathcal{C}_{g,n/R},\bar{x})\xrightarrow{\sim}\pi_{1}^{\prime}(\mathcal{C}_{g,n/R},\bar{y})$, $\bar{\lambda}:\pi_{1}(\mathcal{M}_{g,n/R},\bar{\eta})\xrightarrow{\sim}\pi_{1}(\mathcal{M}_{g,n/R},\bar{\xi})$, and $\tilde{\lambda}:\pi_{1}(C_{\bar{\eta}},\bar{x})^{{(\ell)}}\xrightarrow{\sim}\pi_{1}(C_{\bar{\xi}},\bar{y})^{{(\ell)}}$, which together make the following diagram commute:
(8) ${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(C_{\bar{\xi}},\bar{y})^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}^{\prime}(\mathcal{C}_{g,n/\bar{k}},\bar{y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g,n/\bar{k}},\bar{\xi})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(C_{\bar{\xi}},\bar{y})^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}^{\prime}(\mathcal{C}_{g,n/R},\bar{y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g,n/R},\bar{\xi})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(C_{\bar{\eta}},\bar{x})^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{\lambda}}\scriptstyle{\cong}$${\pi_{1}^{\prime}(\mathcal{C}_{g,n/R},\bar{x})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda^{\prime}}\scriptstyle{\cong}$${\pi_{1}(\mathcal{M}_{g,n/R},\bar{\eta})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\bar{\lambda}}\scriptstyle{\cong}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(C_{\bar{\eta}},\bar{x})^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}^{\prime}(\mathcal{C}_{g,n/\bar{K}},\bar{x})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g,n/\bar{K}},\bar{\eta})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1,}$
where rows are exact (see §2.1). Observe that the monodromy representation $\rho_{\mathcal{C}_{g,n}}:\pi_{1}(\mathcal{C}_{g,n/\bar{k}})\to\operatorname{Sp}(H_{\mathbb{Z}_{\ell}})$ factors through $\pi_{1}(\mathcal{M}_{g,n/\bar{k}})$, and hence through the quotient $\pi_{1}^{\prime}(\mathcal{C}_{g,n/\bar{k}})$. It follows that there is a unique homomorphism $\pi_{1}(\mathcal{C}_{g,n/\bar{k}})^{{\mathrm{rel}(\ell)}}\to\pi_{1}^{\prime}(\mathcal{C}_{g,n/\bar{k}})^{{\mathrm{rel}(\ell)}}$ by naturality of relative pro-$\ell$ completion. On the other hand, we have a commutative diagram
[…]
where $N$ is defined in §2.1. Since $K$ is pro-$\ell$, the image of $N$ in $\pi_{1}(\mathcal{C}_{g,n/\bar{k}})^{{\mathrm{rel}(\ell)}}$ is trivial. Hence the completion $\pi_{1}(\mathcal{C}_{g,n/\bar{k}})\to\pi_{1}(\mathcal{C}_{g,n/\bar{k}})^{{\mathrm{rel}(\ell)}}$ factors through $\pi_{1}^{\prime}(\mathcal{C}_{g,n/\bar{k}})$, inducing a unique map $\pi_{1}^{\prime}(\mathcal{C}_{g,n/\bar{k}})^{{\mathrm{rel}(\ell)}}\to\pi_{1}(\mathcal{C}_{g,n/\bar{k}})^{{\mathrm{rel}(\ell)}}$.
By the universal property of relative pro-$\ell$ completion, these two maps are inverse isomorphisms, and thus $\pi_{1}(\mathcal{C}_{g,n/\bar{k}})^{{\mathrm{rel}(\ell)}}\cong\pi_{1}^{\prime}(\mathcal{C}_{g,n/\bar{k}})^{{\mathrm{rel}(\ell)}}$. The same argument applies over $R$ and $\bar{K}$.
Applying relative pro-$\ell$ completion to the diagram (8), we obtain
(9) ${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(C_{\bar{\xi}},\bar{y})^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{C}_{g,n/\bar{k}},\bar{y})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\pi_{1}(\mathcal{M}_{g,n/\bar{k}},\bar{\xi})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(C_{\bar{\xi}},\bar{y})^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{C}_{g,n/R},\bar{y})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g,n/R},\bar{\xi})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(C_{\bar{\eta}},\bar{x})^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{\lambda}}\scriptstyle{\cong}$${\pi_{1}(\mathcal{C}_{g,n/R},\bar{x})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda^{{}^{\prime}{\mathrm{rel}(\ell)}}}\scriptstyle{\cong}$${\pi_{1}(\mathcal{M}_{g,n/R},\bar{\eta})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\bar{\lambda}^{\mathrm{rel}(\ell)}}\scriptstyle{\cong}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(C_{\bar{\eta}},\bar{x})^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{C}_{g,n/\bar{K}},\bar{x})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\pi_{1}(\mathcal{M}_{g,n/\bar{K}},\bar{\eta})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${1.}$
The rows are exact by right exactness of relative pro-$\ell$ completion and the center-freeness of $\pi_{1}(C_{\bar{\xi}})^{{(\ell)}}(\cong\pi_{1}(C_{\bar{\eta}})^{{(\ell)}})$, and the right vertical homomorphisms are isomorphisms (see the proof of [WAT19, Thm. 5.6]). It therefore follows that all middle vertical homomorphisms are isomorphisms. ∎
4.3. Specialization for configuration spaces in the ordered case
Let $L$ be an algebraically closed field of characteristic 0. Let $\bar{\eta}$ be a geometric point of $\mathcal{M}_{g/\bar{L}}$, and let $C_{\bar{\eta}}$ be the fiber of $\pi_{g/L}:\mathcal{C}_{g/L}\to\mathcal{M}_{g/L}$ over $\bar{\eta}$. Fix a geometric point $\bar{x}\in F(C_{\bar{\eta}},n)$. Recall that $\pi_{1}(F(C_{\bar{\eta}},n),\bar{x})$ is naturally isomorphic to the profinite completion $\widehat{P}_{g,n}$; we identify $\pi_{1}(F(C_{\bar{\eta}},n),\bar{x})$ with $\widehat{P}_{g,n}$.
The forgetful morphism $\mathcal{M}_{g,n/L}\to\mathcal{M}_{g/L}$ yields an exact sequence
$1\longrightarrow\widehat{P}_{g,n}\longrightarrow\pi_{1}(\mathcal{M}_{g,n/L},\bar{x})\longrightarrow\pi_{1}(\mathcal{M}_{g/L},\bar{\eta})\longrightarrow 1.$
Injectivity on the left follows from the center-freeness of $\widehat{P}_{g,n}$, which can be proved using [AND74] and [BIR74, Lem. 4.2.2]. Conjugation in $\pi_{1}(\mathcal{M}_{g,n/L})$ gives a continuous action on $\widehat{P}_{g,n}$, which induces a continuous action on $P_{g,n}^{{(\ell)}}$. Consequently, there is a natural homomorphism
$\rho^{{(\ell)}}:\pi_{1}(\mathcal{M}_{g,n/L})\longrightarrow\operatorname{Aut}_{\mathrm{cts}}(P_{g,n}^{{(\ell)}}).$
Lemma 4.3.
The homomorphism $\rho^{(\ell)}:\pi_{1}({\mathcal{M}}_{g,n/L})\to\operatorname{Aut}_{\mathrm{cts}}(P_{g,n}^{(\ell)})$ factors through the relative pro-$\ell$ completion $\pi_{1}({\mathcal{M}}_{g,n/L})^{{\mathrm{rel}(\ell)}}$.
Proof.
Denote the image of $\rho^{(\ell)}$ by $G\subset\operatorname{Aut}_{\mathrm{cts}}(P_{g,n}^{(\ell)})$. The natural monodromy representation $\pi_{1}({\mathcal{M}}_{g,n/L})\to\operatorname{Sp}(H_{{\mathbb{Z}_{\ell}}})$ factors through $G$, yielding a continuous homomorphism $\psi:G\to\operatorname{Sp}(H_{{\mathbb{Z}_{\ell}}})$.
Since $P_{g,n}$ is finitely generated and $P_{g,n}^{\mathrm{ab}}$ has no $\ell$-primary torsion, the canonical map $P_{g,n}\to P_{g,n}^{{(\ell)}}$ induces
$H_{1}(P_{g,n}^{{(\ell)}},{\mathbb{Z}_{\ell}})\cong H_{1}(P_{g,n},{\mathbb{Z}})\otimes{\mathbb{Z}}_{\ell}.$
Let
$\operatorname{Aut}^{0}_{\mathrm{cts}}(P_{g,n}^{(\ell)}):=\ker\!\Bigl(\operatorname{Aut}_{\mathrm{cts}}(P_{g,n}^{(\ell)})\to\operatorname{Aut}_{\mathrm{cts}}\bigl(H_{1}(P_{g,n})\otimes{\mathbb{Z}}_{\ell}\bigr)\Bigr).$
A standard fact implies that $\operatorname{Aut}^{0}_{\mathrm{cts}}(P_{g,n}^{(\ell)})$ is a pro-$\ell$ group. Moreover, the image of $G$ in $\operatorname{Aut}_{\mathrm{cts}}(H_{1}(P_{g,n})\otimes{\mathbb{Z}_{\ell}})$ is the diagonal copy of $\operatorname{Sp}(H_{{\mathbb{Z}_{\ell}}})$. Thus we have a commutative diagram with exact rows
[…]
In particular, $\ker\psi\subset\operatorname{Aut}^{0}_{\mathrm{cts}}(P_{g,n}^{(\ell)})$ is a pro-$\ell$ group. By the universal property of relative pro-$\ell$ completion, this implies that $\rho^{(\ell)}$ factors through $\pi_{1}({\mathcal{M}}_{g,n/L})^{{\mathrm{rel}(\ell)}}$, i.e. there exists a homomorphism $\pi_{1}({\mathcal{M}}_{g,n/L})^{{\mathrm{rel}(\ell)}}\to G$ whose composition with the inclusion $G\hookrightarrow\operatorname{Aut}_{\mathrm{cts}}(P_{g,n}^{(\ell)})$ equals $\rho^{(\ell)}$. ∎
Recall the notation from §3.1. Now return to the mixed-characteristic model $X/R$ used to define specialization. Let $C$ be a smooth proper curve over $\bar{k}$, and let $X\to\operatorname{Spec}R$ be a smooth proper $R$-scheme whose special fiber is isomorphic to $C$. There is a unique morphism $\psi:\operatorname{Spec}R\to\mathcal{M}_{g/R}$ such that the pullback of the universal curve $\mathcal{C}_{g}\to\mathcal{M}_{g}$ along $\psi$ is isomorphic to $X\to\operatorname{Spec}R$; we denote this pullback by $X_{R}$. Let $\phi:X_{R}\to\mathcal{C}_{g/R}$ be the induced morphism.
The morphism $\phi:X_{R}\to\mathcal{C}_{g/R}$ is a monomorphism; hence the induced morphism $\phi^{n}:X_{R}^{n}\to\mathcal{C}_{g/R}^{n}$ preserves the complement of the fat diagonal and therefore induces a morphism
$F(\phi):F(X_{R},n)\longrightarrow F(\mathcal{C}_{g/R},n).$
Since $\mathcal{M}_{g,n/R}$ is canonically isomorphic to $F(\mathcal{C}_{g/R},n)$, we obtain a sequence of morphisms
(10) $F(X_{R},n)\xrightarrow{\,F(\phi)\,}\mathcal{M}_{g,n/R}\longrightarrow\mathcal{M}_{g/R}.$
Let $\bar{\xi}$ and $\bar{\eta}$ be geometric points of $\mathcal{M}_{g/\bar{k}}$ and $\mathcal{M}_{g/\bar{K}}$, respectively. Via the structure morphism $\mathcal{M}_{g/R}\to\operatorname{Spec}R$, we also regard $\bar{\xi}$ and $\bar{\eta}$ as the closed geometric point and the generic geometric point of $\operatorname{Spec}R$, respectively.
Over $\bar{\eta}$, we have $F(X_{R},n)_{\bar{\eta}}\cong F(X_{\bar{\eta}},n)$, and there is an exact sequence of étale fundamental groups
$1\longrightarrow\pi_{1}\bigl(F(X_{\bar{\eta}},n),\bar{x}\bigr)\longrightarrow\pi_{1}\bigl(\mathcal{M}_{g,n/\bar{K}},\bar{x}\bigr)\longrightarrow\pi_{1}\bigl(\mathcal{M}_{g/\bar{K}},\bar{\eta}\bigr)\longrightarrow 1.$
On the other hand, over the geometric point $\bar{\xi}$, we have $F(X_{R},n)_{\bar{\xi}}\cong F(C,n)$, and there is a sequence
$\pi_{1}\bigl(F(C,n),\bar{y}\bigr)\longrightarrow\pi_{1}\bigl(\mathcal{M}_{g,n/\bar{k}},\bar{y}\bigr)\longrightarrow\pi_{1}\bigl(\mathcal{M}_{g/\bar{k}},\bar{\xi}\bigr)\longrightarrow 1,$
where the composition is trivial and the second homomorphism is surjective.
Proposition 4.4.
For $g\geq 2$ and $n\geq 1$, there is a commutative diagram:
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(F(C,n),\bar{y})^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g,n/\bar{k}},\bar{y})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g/\bar{k}},\bar{\xi})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${P^{{(\ell)}}_{g,n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\pi_{1}(\mathcal{M}_{g,n/\bar{K}},\bar{x})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\pi_{1}(\mathcal{M}_{g/\bar{K}},\bar{\eta})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${1,}$
where the rows are exact and all vertical maps are isomorphisms.
Proof.
Fix the base-point change isomorphism
$\lambda:\pi_{1}\bigl(F(X_{R},n),\bar{x}\bigr)\xrightarrow{\cong}\pi_{1}\bigl(F(X_{R},n),\bar{y}\bigr),$
where $\bar{x}$ (resp. $\bar{y}$) lies over $\bar{\eta}$ (resp. $\bar{\xi}$). Via the sequence (10), the isomorphism $\lambda$ induces isomorphisms
$\lambda^{\prime}:\pi_{1}\bigl(\mathcal{M}_{g,n/R},\bar{x}\bigr)\xrightarrow{\cong}\pi_{1}\bigl(\mathcal{M}_{g,n/R},\bar{y}\bigr),\qquad\lambda^{\prime\prime}:\pi_{1}\bigl(\mathcal{M}_{g/R},\bar{\eta}\bigr)\xrightarrow{\cong}\pi_{1}\bigl(\mathcal{M}_{g/R},\bar{\xi}\bigr).$
These fit into a commutative diagram
${\pi_{1}(F(X_{R},n),\bar{x})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F(\phi)_{*}}$$\scriptstyle{\lambda}\scriptstyle{\cong}$${\pi_{1}(\mathcal{M}_{g,n/R},\bar{x})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda^{\prime}}\scriptstyle{\cong}$${\pi_{1}(\mathcal{M}_{g/R},\bar{\eta})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda^{\prime\prime}}\scriptstyle{\cong}$${\pi_{1}(F(X_{R},n),\bar{y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F(\phi)_{*}}$${\pi_{1}(\mathcal{M}_{g,n/R},\bar{y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g/R},\bar{\xi}).}$
Passing to fibers over the generic and special geometric points yields the following commutative diagram, whose top row is exact:
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(F(X_{\bar{\eta}},n),\bar{x})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g,n/\bar{K}},\bar{x})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g/\bar{K}},\bar{\eta})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${\pi_{1}(F(X_{R},n),\bar{x})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda}\scriptstyle{\cong}$${\pi_{1}(\mathcal{M}_{g,n/R},\bar{x})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda^{\prime}}\scriptstyle{\cong}$${\pi_{1}(\mathcal{M}_{g/R},\bar{\eta})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda^{\prime\prime}}\scriptstyle{\cong}$${1}$${\pi_{1}(F(X_{R},n),\bar{y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g,n/R},\bar{y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g/R},\bar{\xi})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${\pi_{1}(F(C,n),\bar{y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g,n/\bar{k}},\bar{y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g/\bar{k}},\bar{\xi})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
Applying relative pro-$\ell$ completion, we obtain a commutative diagram
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(F(X_{\bar{\eta}},n),\bar{x})^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\pi_{1}(\mathcal{M}_{g,n/\bar{K}},\bar{x})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\pi_{1}(\mathcal{M}_{g/\bar{K}},\bar{\eta})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(F(X_{R},n),\bar{x})^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda}\scriptstyle{\cong}$${\pi_{1}(\mathcal{M}_{g,n/R},\bar{x})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda^{\prime}}\scriptstyle{\cong}$${\pi_{1}(\mathcal{M}_{g/R},\bar{\eta})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda^{\prime\prime}}\scriptstyle{\cong}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(F(X_{R},n),\bar{y})^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g,n/R},\bar{y})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g/R},\bar{\xi})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(F(C,n),\bar{y})^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\pi_{1}(\mathcal{M}_{g,n/\bar{k}},\bar{y})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\pi_{1}(\mathcal{M}_{g/\bar{k}},\bar{\xi})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${1.}$
The left vertical maps are the specialization isomorphisms defined in §3.1. The middle and right vertical maps are isomorphisms by the proof of [WAT19, Thm. 5.6].
It remains to justify exactness of the top row. By Lemma 4.3, the conjugation action of $\pi_{1}(\mathcal{M}_{g,n/\bar{K}})$ on $\pi_{1}(F(X_{\bar{\eta}},n))$ induces an action of $\pi_{1}(\mathcal{M}_{g,n/\bar{K}})^{{\mathrm{rel}(\ell)}}$ on $\pi_{1}(F(X_{\bar{\eta}},n))^{{(\ell)}}$. Let $W_{\bullet}$ denote the weight filtration on $\pi_{1}(F(X_{\bar{\eta}},n))^{{(\ell)}}$ constructed in [NTU95]. Since the associated graded Lie algebra $\operatorname{Gr}^{W}_{\bullet}\pi_{1}(F(X_{\bar{\eta}},n))^{{(\ell)}}$ has trivial center and the filtration $W_{\bullet}$ is separated ([NTU95, Cor. 2.6, p. 201]), it follows that $\pi_{1}(F(X_{\bar{\eta}},n))^{{(\ell)}}$ is center-free. Consequently, the induced homomorphism
$\pi_{1}(F(X_{\bar{\eta}},n))^{{(\ell)}}\longrightarrow\pi_{1}(\mathcal{M}_{g,n/\bar{K}})^{{\mathrm{rel}(\ell)}}$
is injective. This proves exactness of the top row, and hence all rows in the diagram are exact. The proposition follows. ∎
4.4. Homotopy exact sequences in the unordered case
We first compare the ordered and unordered universal curves. Consider the commutative diagram
${{\mathcal{C}}_{g,n/\bar{k}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{M}}_{g,n/\bar{k}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{C}}_{g,[n]/\bar{k}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{M}}_{g,[n]/\bar{k}},}$
which induces a commutative diagram of fundamental groups
(11) ${\pi_{1}\!\bigl(C,\bar{y}\bigr)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}({\mathcal{C}}_{g,n/\bar{k}},\bar{y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}({\mathcal{M}}_{g,n/\bar{k}},\bar{\xi})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${\pi_{1}\!\bigl(C,\bar{y}\bigr)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}({\mathcal{C}}_{g,[n]/\bar{k}},\bar{y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}({\mathcal{M}}_{g,[n]/\bar{k}},\bar{\xi})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1,}$
where the rows are exact on the right.
Since the $S_{n}$-action permutes the marked points but acts trivially on the underlying curve, the local system $\mathbb{H}_{\mathbb{Z}_{\ell}}$ descends to ${\mathcal{M}}_{g,[n]}$. Consequently, the representation
$\rho_{\mathbb{Z}_{\ell},\bar{\xi}}:\pi_{1}({\mathcal{M}}_{g,n/\bar{k}},\bar{\xi})\longrightarrow\operatorname{Sp}(H_{\mathbb{Z}_{\ell}})$
factors through $\rho_{{\mathbb{Z}_{\ell}},[\bar{\xi}]}:\pi_{1}({\mathcal{M}}_{g,[n]/\bar{k}},\bar{\xi})\to\operatorname{Sp}(H_{\mathbb{Z}_{\ell}})$. Thus, there is a commutative diagram
(12) ${\pi_{1}({\mathcal{C}}_{g,n/\bar{k}},\bar{y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}({\mathcal{M}}_{g,n/\bar{k}},\bar{\xi})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rho_{{\mathbb{Z}_{\ell}},\bar{\xi}}}$${\operatorname{Sp}(H_{\mathbb{Z}_{\ell}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}({\mathcal{C}}_{g,[n]/\bar{k}},\bar{y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}({\mathcal{M}}_{g,[n]/\bar{k}},\bar{\xi})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rho_{{\mathbb{Z}_{\ell}},[\bar{\xi}]}}$${\operatorname{Sp}(H_{\mathbb{Z}_{\ell}}).}$
Since $\pi_{1}(C)^{{(\ell)}}$ is center-free, applying relative pro-$\ell$ completion to (12) with respect to the representations into $\operatorname{Sp}(H_{\mathbb{Z}_{\ell}})$ yields the following result.
Proposition 4.5.
For $g\geq 2$ and $n\geq 0$, the diagram (11) induces a commutative diagram
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(C)^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}({\mathcal{C}}_{g,n/\bar{k}})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}({\mathcal{M}}_{g,n/\bar{k}})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(C)^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}({\mathcal{C}}_{g,[n]/\bar{k}})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}({\mathcal{M}}_{g,[n]/\bar{k}})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1,}$
whose rows are exact.
We next compare the ordered and unordered configuration spaces. Consider the commutative diagram
${\mathcal{M}_{g,n/\bar{k}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}_{g/\bar{k}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}_{g,[n]/\bar{k}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}_{g/\bar{k}},}$
which induces a commutative diagram of fundamental groups
(13) ${\pi_{1}\!\bigl(F(C,n),\bar{y}\bigr)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g,n/\bar{k}},\bar{y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g/\bar{k}},\bar{\xi})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${\pi_{1}\!\bigl(F(C,[n]),\bar{y}\bigr)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g,[n]/\bar{k}},\bar{y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g/\bar{k}},\bar{\xi})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1,}$
whose rows are exact on the right. In what follows, we suppress basepoint notation for fundamental groups whenever it is clear from the context.
Proposition 4.6.
Suppose $g\geq 2$ and $n\geq 1$. Let $\ell\neq p$ be an odd prime. There is a commutative diagram
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}\!\bigl(F(C,n)\bigr)^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g,n/\bar{k}})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g/\bar{k}})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${\pi_{1}\!\bigl(F(C,[n])\bigr)^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g,[n]/\bar{k}})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g/\bar{k}})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1,}$
whose rows are exact.
Proof.
Consider the exact sequence
$1\longrightarrow\pi_{1}({\mathcal{M}}_{g,n/\bar{k}})\longrightarrow\pi_{1}({\mathcal{M}}_{g,[n]/\bar{k}})\longrightarrow S_{n}\longrightarrow 1,$
which fits into the commutative diagram
(14) ${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}({\mathcal{M}}_{g,n/\bar{k}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rho_{\mathbb{Z}_{\ell},\bar{\xi}}}$${\pi_{1}({\mathcal{M}}_{g,[n]/\bar{k}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rho_{\mathbb{Z}_{\ell},[\bar{\xi}]}}$${S_{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Sp}(H_{\mathbb{Z}_{\ell}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Sp}(H_{\mathbb{Z}_{\ell}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$
For $n\geq 2$ and $\ell$ an odd prime, the pro-$\ell$ completion of $S_{n}$ is trivial. Since relative pro-$\ell$ completion is right exact [HM09, Prop. 2.4], applying it to (14) yields an exact sequence
$\pi_{1}({\mathcal{M}}_{g,n/\bar{k}})^{{\mathrm{rel}(\ell)}}\longrightarrow\pi_{1}({\mathcal{M}}_{g,[n]/\bar{k}})^{{\mathrm{rel}(\ell)}}\longrightarrow S_{n}^{{(\ell)}}\cong 1.$
Hence the induced map $\pi_{1}({\mathcal{M}}_{g,n/\bar{k}})^{{\mathrm{rel}(\ell)}}\to\pi_{1}({\mathcal{M}}_{g,[n]/\bar{k}})^{{\mathrm{rel}(\ell)}}$ is surjective.
Applying relative pro-$\ell$ completion to (13), we obtain a commutative diagram
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}\!\bigl(F(C,n)\bigr)^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g,n/\bar{k}})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g/\bar{k}})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${\pi_{1}\!\bigl(F(C,[n])\bigr)^{{(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g,[n]/\bar{k}})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathcal{M}_{g/\bar{k}})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
The first row is exact by Proposition 4.4. By right exactness of the completion, the second row is exact on the right. Since the middle vertical map is surjective, the second row is exact in the middle as well. ∎
5. Continuous Relative Completion
The continuous relative completion is the key tool used in the prequel to this paper [LW25]. It is the profinite analogue of the relative completion of a discrete group. Detailed introductions to the definition and properties of relative completion can be found in [HAI93, HAI98, HAI08].
[…]
$1\longrightarrow{\mathcal{U}}\longrightarrow{\mathcal{G}}\longrightarrow R\longrightarrow 1,$
[…]
commutes.
5.1. Continuous relative completion of $\pi_{1}({\mathcal{C}}_{g,n/\bar{\mathbb{Q}}})\to\pi_{1}({\mathcal{M}}_{g,n/\bar{\mathbb{Q}}})$
We briefly recall the structure of the continuous relative completions of $\pi_{1}({\mathcal{M}}_{g,n/\bar{\mathbb{Q}}})$ and $\pi_{1}({\mathcal{C}}_{g,n/\bar{\mathbb{Q}}})$ associated with the universal curve, following the framework of Hain and Matsumoto.
Consider the $\ell$-adic local system
$\mathbb{H}_{{\mathbb{Q}}_{\ell}}:=\mathbb{H}$$_{\mathbb{Z}_{\ell}}\otimes{\mathbb{Q}}_{\ell}$
on $\mathcal{M}_{g,n/\mathbb{Z}}$. Let $\bar{\zeta}$ be a geometric point of ${\mathcal{M}}_{g,n/\bar{\mathbb{Q}}}$, and let $\bar{a}$ be a geometric point of ${\mathcal{C}}_{g,n/\bar{\mathbb{Q}}}$ lying over $\bar{\zeta}$. Write $H_{{\mathbb{Q}}_{\ell}}$ for the fiber of $\mathbb{H}_{{\mathbb{Q}}_{\ell}}$ over $\bar{\zeta}$.
The associated monodromy representation
$\rho_{\bar{\zeta}}:\pi_{1}({\mathcal{M}}_{g,n/\bar{\mathbb{Q}}},\bar{\zeta})\longrightarrow\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell}})$
is continuous with Zariski-dense image. Composing $\rho_{\bar{\zeta}}$ with the homomorphism induced by the universal curve, $\pi_{1}({\mathcal{C}}_{g,n/\bar{\mathbb{Q}}},\bar{a})\longrightarrow\pi_{1}({\mathcal{M}}_{g,n/\bar{\mathbb{Q}}},\bar{\zeta})$, we obtain a representation
$\rho_{{\mathcal{C}},\bar{\zeta}}:\pi_{1}({\mathcal{C}}_{g,n/\bar{\mathbb{Q}}},\bar{a})\longrightarrow\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell}}),$
which corresponds to the pullback of the local system $\mathbb{H}_{{\mathbb{Q}}_{\ell}}$ along the universal curve.
Denote by ${\mathcal{G}}_{g,n}$ and ${\mathcal{G}}_{{\mathcal{C}}_{g,n}}$ the continuous relative completions of $\pi_{1}({\mathcal{M}}_{g,n/\bar{\mathbb{Q}}})$ and $\pi_{1}({\mathcal{C}}_{g,n/\bar{\mathbb{Q}}})$ with respect to $\rho_{\bar{\zeta}}$ and $\rho_{{\mathcal{C}},\bar{\zeta}}$, respectively. Let ${\mathcal{U}}_{g,n}$ and ${\mathcal{U}}_{{\mathcal{C}}_{g,n}}$ be their unipotent radicals. We write
${\mathfrak{g}}_{g,n},\ {\mathfrak{g}}_{{\mathcal{C}}_{g,n}},\ {\mathfrak{u}}_{g,n},\ {\mathfrak{u}}_{{\mathcal{C}}_{g,n}}$
for the Lie algebras of ${\mathcal{G}}_{g,n}$, ${\mathcal{G}}_{{\mathcal{C}}_{g,n}}$, ${\mathcal{U}}_{g,n}$, and ${\mathcal{U}}_{{\mathcal{C}}_{g,n}}$, respectively.
These Lie algebras carry natural weight filtrations, inherited from the corresponding relative completions over $\mathbb{Q}$} constructed using Hodge-theoretic methods. We use only the induced structural properties of these filtrations (cf. [HAI98, Thm. 13.1]).
5.2. The relative completion and relative pro-$\ell$ completion
Suppose that $\Gamma$ is a profinite group and let
$\rho_{\mathbb{Z}_{\ell}}:\Gamma\longrightarrow R(\mathbb{Z}_{\ell})$
be a continuous homomorphism whose extension
$\rho_{\mathbb{Q}_{\ell}}:\Gamma\longrightarrow R(\mathbb{Q}_{\ell})$
has Zariski–dense image. Let ${\Gamma}^{\mathrm{rel}(\ell)}$ denote the relative pro-$\ell$ completion of $\Gamma$ with respect to $\rho_{\mathbb{Z}_{\ell}}$, and write
$\rho^{{\mathrm{rel}(\ell)}}_{\mathbb{Z}_{\ell}}:\Gamma^{{\mathrm{rel}(\ell)}}\longrightarrow R(\mathbb{Z}_{\ell})$
for the induced homomorphism. Since $\rho(\Gamma)\subseteq\rho^{{\mathrm{rel}(\ell)}}_{\mathbb{Z}_{\ell}}(\Gamma^{{\mathrm{rel}(\ell)}})$, the induced map
$\rho^{\mathrm{rel}(\ell)}_{{{\mathbb{Q}}_{\ell}}}:\Gamma^{{\mathrm{rel}(\ell)}}\longrightarrow R(\mathbb{Q}_{\ell})$
also has Zariski–dense image.
The following observation is then an immediate consequence of the universal property.
Proposition 5.1.
The continuous relative completion of $\Gamma^{{\mathrm{rel}(\ell)}}$ with respect to the representation $\rho^{{\mathrm{rel}(\ell)}}_{{\mathbb{Q}}_{\ell}}$ is canonically isomorphic to the continuous relative completion of $\Gamma$ with respect to $\rho_{{\mathbb{Q}}_{\ell}}$.
Proof.
Let $\mathcal{G}$} and ${\mathcal{G}}^{{\mathrm{rel}(\ell)}}$ be the continuous relative completions of $\Gamma$} and ${\Gamma}^{{\mathrm{rel}(\ell)}}$ with respect to $\rho_{{\mathbb{Q}}_{\ell}}$ and $\rho^{{\mathrm{rel}(\ell)}}_{{\mathbb{Q}}_{\ell}}$, respectively. The surjection ${\Gamma}\to{\Gamma}^{{\mathrm{rel}(\ell)}}$ induces a unique surjective morphism $\phi:{\mathcal{G}}\to{\mathcal{G}}^{{\mathrm{rel}(\ell)}}$. Let $\bar{\Gamma}$ be the image of $\Gamma$} in $\mathcal{G}$}. Then $\bar{\Gamma}$ is an extension of the image of $\rho_{{\mathbb{Z}_{\ell}}}$ by a pro-$\ell$ group:
[…]
Since every compact subgroup of $U({\mathbb{Q}}_{\ell})$ is pro-$\ell$, the kernel $\ker(\rho_{\mathbb{Z}_{\ell}})$ is a pro-$\ell$ group. Hence the map ${\Gamma}\to\bar{{\Gamma}}$ factors through ${\Gamma}^{{\mathrm{rel}(\ell)}}$. Since the diagram
${{\Gamma}^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rho^{{\mathrm{rel}(\ell)}}_{{\mathbb{Q}}_{\ell}}}$${{\mathcal{G}}({{\mathbb{Q}}_{\ell}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R({{\mathbb{Q}}_{\ell}})}$
commutes, there exists a unique surjective morphism $\psi:{\mathcal{G}}^{{\mathrm{rel}(\ell)}}\to{\mathcal{G}}$. By the universal property of relative completion, $\phi$ and $\psi$ are mutually inverse, and therefore give a canonical isomorphism. ∎
5.3. Continuous relative completion in the ordered case
Consider the $\ell$-adic local system $\mathbb{H}_{{\mathbb{Q}}_{\ell}}:=\mathbb{H}$$_{\mathbb{Z}_{\ell}}\otimes{\mathbb{Q}}_{\ell}$ on $\mathcal{M}_{g,n/\mathbb{Z}}$. With the notation of §4.2, write $H_{{\mathbb{Q}}_{\ell},\bar{\xi}}$ (resp. $H_{{\mathbb{Q}}_{\ell},\bar{\eta}}$) for its fiber over $\bar{\xi}$ (resp. $\bar{\eta}$). This yields symplectic representations
$\rho_{{\mathbb{Q}}_{\ell},\bar{\xi}}:\pi_{1}(\mathcal{M}_{g,n/\bar{k}},\bar{\xi})\longrightarrow\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell},\bar{\xi}}),\qquad\rho_{{\mathbb{Q}}_{\ell},\bar{\eta}}:\pi_{1}(\mathcal{M}_{g,n/\bar{K}},\bar{\eta})\longrightarrow\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell},\bar{\eta}}).$
Passing to relative pro-$\ell$ completions, we obtain the induced representations
$\rho^{{\mathrm{rel}(\ell)}}_{{\mathbb{Q}}_{\ell},\bar{\xi}}:\pi_{1}(\mathcal{M}_{g,n/\bar{k}},\bar{\xi})^{{\mathrm{rel}(\ell)}}\longrightarrow\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell},\bar{\xi}}),\quad\rho^{{\mathrm{rel}(\ell)}}_{{\mathbb{Q}}_{\ell},\bar{\eta}}:\pi_{1}(\mathcal{M}_{g,n/\bar{K}},\bar{\eta})^{{\mathrm{rel}(\ell)}}\longrightarrow\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell},\bar{\eta}}).$
Denote by ${\mathcal{G}}_{g,n,\bar{\xi}}$ (resp. ${\mathcal{G}}_{g,n,\bar{\eta}}$) the continuous relative completion of $\pi_{1}({\mathcal{M}}_{g,n/\bar{k}},\bar{\xi})$ (resp. $\pi_{1}({\mathcal{M}}_{g,n/\bar{K}},\bar{\eta})$) with respect to $\rho_{{\mathbb{Q}}_{\ell},\bar{\xi}}$ (resp. $\rho_{{\mathbb{Q}}_{\ell},\bar{\eta}}$) Similarly, denote by ${\mathcal{G}}_{{\mathcal{C}}_{g,n},\bar{\xi}}$ (resp. ${\mathcal{G}}_{{\mathcal{C}}_{g,n},\bar{\eta}}$) the continuous relative completion of $\pi_{1}({\mathcal{C}}_{g,n/\bar{k}},\bar{\xi})$ (resp. $\pi_{1}({\mathcal{C}}_{g,n/\bar{K}},\bar{\eta})$) with respect to the induced representations into $\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell},\bar{\xi}})$ (resp. $\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell},\bar{\eta}})$).
Proposition 5.2.
[…]
${\mathcal{G}}_{g,n,\bar{\eta}}\;\cong\;{\mathcal{G}}_{g,n,\bar{\xi}}\;\cong\;{\mathcal{G}}_{g,n}$
and
${\mathcal{G}}_{{\mathcal{C}}_{g,n},\bar{\eta}}\cong{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathcal{G}}_{{\mathcal{C}}_{g,n},\bar{\xi}}}\cong{\mathcal{G}}_{{\mathcal{C}}_{g,n}}.$
Proof.
By Proposition 4.2 and Lemma 3.1, the representations $\rho^{{\mathrm{rel}(\ell)}}_{{\mathbb{Q}}_{\ell},\bar{\xi}}$ and $\rho^{{\mathrm{rel}(\ell)}}_{{\mathbb{Q}}_{\ell},\bar{\eta}}$, when composed with the natural morphisms
$\pi_{1}({\mathcal{C}}_{g,n/\bar{k}})^{{\mathrm{rel}(\ell)}}\to\pi_{1}({\mathcal{M}}_{g,n/\bar{k}})^{{\mathrm{rel}(\ell)}},\qquad\pi_{1}({\mathcal{C}}_{g,n/\bar{K}})^{{\mathrm{rel}(\ell)}}\to\pi_{1}({\mathcal{M}}_{g,n/\bar{K}})^{{\mathrm{rel}(\ell)}},$
induce, on $H_{1}$, the actions arising from conjugation on $\pi_{1}(C_{\bar{\xi}})^{{(\ell)}}$ and $\pi_{1}(C_{\bar{\eta}})^{{(\ell)}}$, respectively. Consequently, the specialization isomorphism yields a commutative diagram
${\pi_{1}({\mathcal{C}}_{g,n/\bar{K}},\bar{\eta})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\pi_{1}({\mathcal{C}}_{g,n/\bar{k}},\bar{\xi})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}({\mathcal{M}}_{g,n/\bar{K}},\bar{\eta})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{\rho^{{\mathrm{rel}(\ell)}}_{{\mathbb{Q}}_{\ell},\bar{\eta}}}$${\pi_{1}({\mathcal{M}}_{g,n/\bar{k}},\bar{\xi})^{{\mathrm{rel}(\ell)}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rho^{{\mathrm{rel}(\ell)}}_{{\mathbb{Q}}_{\ell},\bar{\xi}}}$${\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell},\bar{\eta}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell},\bar{\xi}}),}$
where the bottom isomorphism is induced by $\pi_{1}(C_{\bar{\eta}})^{{(\ell)}}\cong\pi_{1}(C_{\bar{\xi}})^{{(\ell)}}$. Taking the continuous relative completion of the above diagram and using Proposition 5.1, which identifies the resulting completion with the continuous relative completion taken with respect to the original representations $\rho_{{\mathbb{Q}}_{\ell},\bar{\eta}}$ and $\rho_{{\mathbb{Q}}_{\ell},\bar{\xi}}$, we obtain natural isomorphisms
${\mathcal{G}}_{g,n,\bar{\eta}}\;\cong\;{\mathcal{G}}_{g,n,\bar{\xi}}\quad\text{and}\quad{\mathcal{G}}_{{\mathcal{C}}_{g,n},\bar{\eta}}\;\cong\;{\mathcal{G}}_{{\mathcal{C}}_{g,n},\bar{\xi}}.$
After fixing an embedding $\overline{{\mathbb{Q}}}\hookrightarrow\bar{K}$, invariance of the étale fundamental group under extension of algebraically closed fields of characteristic 0 gives canonical isomorphisms
$\pi_{1}({\mathcal{M}}_{g,n/\bar{K}})\;\xrightarrow{\;\sim\;}\;\pi_{1}({\mathcal{M}}_{g,n/\overline{{\mathbb{Q}}}})\quad\text{and}\quad\pi_{1}({\mathcal{C}}_{g,n/\bar{K}})\;\xrightarrow{\;\sim\;}\;\pi_{1}({\mathcal{C}}_{g,n/\overline{{\mathbb{Q}}}}).$
Applying continuous relative completion, we obtain canonical isomorphisms
${\mathcal{G}}_{g,n,\bar{\eta}}\;\cong\;{\mathcal{G}}_{g,n}\quad\text{and}\quad{\mathcal{G}}_{{\mathcal{C}}_{g,n},\bar{\eta}}\;\cong\;{\mathcal{G}}_{{\mathcal{C}}_{g,n}}.$
5.4. Continuous relative completion in the unordered case
We retain the notation from the ordered case and adapt it to the unordered setting. For simplicity, set $H_{{\mathbb{Q}}_{\ell}}:=H_{{\mathbb{Q}}_{\ell},\bar{\xi}}$. The $\ell$-adic representation $\rho_{{\mathbb{Q}}_{\ell},\bar{\xi}}:\pi_{1}({\mathcal{M}}_{g,n/\bar{k}},\bar{\xi})\longrightarrow\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell}})$ factors through the natural homomorphism $\pi_{1}({\mathcal{M}}_{g,n/\bar{k}})\longrightarrow\pi_{1}({\mathcal{M}}_{g,[n]/\bar{k}})$, and hence induces a representation
$\rho_{{\mathbb{Q}}_{\ell},[\bar{\xi}]}:\pi_{1}({\mathcal{M}}_{g,[n]/\bar{k}},\bar{\xi})\longrightarrow\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell}}).$
Let
$\rho_{{\mathcal{C}}_{{\mathbb{Q}}_{\ell}},[\bar{\xi}]}:\pi_{1}({\mathcal{C}}_{g,[n]/\bar{k}},\bar{x})\longrightarrow\pi_{1}({\mathcal{M}}_{g,[n]/\bar{k}},\bar{\xi})\longrightarrow\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell}})$
denote the induced representation obtained by pullback along the universal curve.
Denote by ${\mathcal{G}}_{g,[n],\bar{\xi}}$ and ${\mathcal{G}}_{{\mathcal{C}}_{g,[n]},\bar{\xi}}$ the continuous relative completions of $\pi_{1}({\mathcal{M}}_{g,[n]/\bar{k}},\bar{\xi})$ and $\pi_{1}({\mathcal{C}}_{g,[n]/\bar{k}},\bar{x})$ with respect to $\rho_{{\mathbb{Q}}_{\ell},[\bar{\xi}]}$ and $\rho_{{\mathcal{C}}_{{\mathbb{Q}}_{\ell}},[\bar{\xi}]}$, respectively. Let ${\mathcal{U}}_{g,[n],\bar{\xi}}$ and ${\mathcal{U}}_{{\mathcal{C}}_{g,[n]},\bar{\xi}}$ be their unipotent radicals, and write
${\mathfrak{g}}_{g,[n],\bar{\xi}},\ {\mathfrak{g}}_{{\mathcal{C}}_{g,[n]},\bar{\xi}},\ {\mathfrak{u}}_{g,[n],\bar{\xi}},\ {\mathfrak{u}}_{{\mathcal{C}}_{g,[n]},\bar{\xi}}$
for the Lie algebras of ${\mathcal{G}}_{g,[n],\bar{\xi}}$, ${\mathcal{G}}_{{\mathcal{C}}_{g,[n]},\bar{\xi}}$, ${\mathcal{U}}_{g,[n],\bar{\xi}}$, and ${\mathcal{U}}_{{\mathcal{C}}_{g,[n]},\bar{\xi}}$, respectively.
Unlike the ordered case, we do not know whether the continuous relative completions in characteristic 0 and characteristic $p$ admit a direct comparison in the unordered setting. Nevertheless, the following result provides the structural input needed for our later arguments and may be viewed as the characteristic $p$ analogue of [LW25, Lem. 5.22].
Lemma 5.3.
For $g\geq 2$ and $n\geq 1$, there is an isomorphism
$\operatorname{Hom}_{\operatorname{Sp}}^{\mathrm{cts}}(H_{1}({\mathfrak{u}}_{g,[n],\bar{\xi}}),H_{{\mathbb{Q}}_{\ell}})\cong{\mathbb{Q}}_{\ell}.$
Proof.
By the cohomological description of relative completion [HAI08, Thm. 3.8], there is a natural isomorphism
$\operatorname{Hom}_{\operatorname{Sp}}^{\mathrm{cts}}\!\bigl(H_{1}({\mathfrak{u}}_{g,[n],\bar{\xi}}),H_{{\mathbb{Q}}_{\ell}}\bigr)\;\cong\;H^{1}_{\mathrm{cts}}\!\bigl(\pi_{1}({\mathcal{M}}_{g,[n]/\bar{k}}),H_{{\mathbb{Q}}_{\ell}}\bigr).$
Since the monodromy representation $\pi_{1}({\mathcal{M}}_{g,n/\bar{k}})\to\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell}})$ is Zariski dense, we have $H_{{\mathbb{Q}}_{\ell}}^{\pi_{1}({\mathcal{M}}_{g,n/\bar{k}})}=0$. Applying the Hochschild–Serre five-term exact sequence associated with $1\to\pi_{1}({\mathcal{M}}_{g,n/\bar{k}})\to\pi_{1}({\mathcal{M}}_{g,[n]/\bar{k}})\to S_{n}\to 1$ therefore yields
$H^{1}_{\mathrm{cts}}\!\bigl(\pi_{1}({\mathcal{M}}_{g,[n]/\bar{k}}),H_{{\mathbb{Q}}_{\ell}}\bigr)\cong H^{1}_{\mathrm{cts}}\!\bigl(\pi_{1}({\mathcal{M}}_{g,n/\bar{k}}),H_{{\mathbb{Q}}_{\ell}}\bigr)^{S_{n}}.$
Again using [HAI08, Thm. 3.8], we have natural isomorphisms
$H^{1}_{\mathrm{cts}}\!\bigl(\pi_{1}({\mathcal{M}}_{g,n/\bar{k}}),H_{{\mathbb{Q}}_{\ell}}\bigr)$ $\cong\operatorname{Hom}_{\operatorname{Sp}}^{\mathrm{cts}}\!\bigl(H_{1}({\mathfrak{u}}_{g,n,\bar{\xi}}),H_{{\mathbb{Q}}_{\ell}}\bigr)$
$\cong\operatorname{Hom}_{\operatorname{Sp}}^{\mathrm{cts}}\!\bigl(H_{1}({\mathfrak{u}}_{g,n,\bar{\eta}}),H_{{\mathbb{Q}}_{\ell},\bar{\eta}}\bigr)$
$\cong H^{1}_{\mathrm{cts}}\!\bigl(\pi_{1}({\mathcal{M}}_{g,n/\bar{K}},\bar{\eta}),H_{{\mathbb{Q}}_{\ell},\bar{\eta}}\bigr)$
$\cong\bigoplus_{j=1}^{n}{\mathbb{Q}}_{\ell},$
where the second isomorphism is induced by Proposition 5.2, and the final isomorphism is induced by the characteristic classes associated with the tautological sections of the universal curve (cf. [LW25, §5.6–§5.7]).
The symmetric group $S_{n}$ acts by permuting the summands of $\bigoplus_{j=1}^{n}{\mathbb{Q}}_{\ell}$. Taking invariants therefore yields
$H^{1}_{\mathrm{cts}}\!\bigl(\pi_{1}({\mathcal{M}}_{g,[n]/\bar{k}}),H_{{\mathbb{Q}}_{\ell}}\bigr)\;\cong\;{\mathbb{Q}}_{\ell},$
which proves the result. ∎
5.5. Key exact sequences in relative completion
We now introduce the exact sequences in continuous relative completion that play a key role in the proof of the main results.
Denote by ${\mathfrak{p}}^{\mathrm{an}}_{g,n}$ the Lie algebra of the $\ell$-adic unipotent completion of $P_{g,n}$. In the case of $n=1$, denote ${\mathfrak{p}}^{\mathrm{an}}_{g,1}$ by ${\mathfrak{p}}^{\mathrm{an}}_{g}$.
5.5.1. Key exact sequences for the universal curve
[…]
(15) ${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F^{\mathrm{un}}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{G}}_{{\mathcal{C}}_{g,n},\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{G}}_{g,n,\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F^{\mathrm{un}}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{G}}_{{\mathcal{C}}_{g,[n]},\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{G}}_{g,[n],\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$
[…]
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F^{\mathrm{un}}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{U}}_{{\mathcal{C}}_{g,n},\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{U}}_{g,n,\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F^{\mathrm{un}}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{U}}_{{\mathcal{C}}_{g,[n]},\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{U}}_{g,[n],\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
Passing to Lie algebras, we obtain a commutative diagram of pronilpotent Lie algebras with exact rows:
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathfrak{p}}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathfrak{u}}_{{\mathcal{C}}_{g,n},\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathfrak{u}}_{g,n,\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathfrak{p}}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathfrak{u}}_{{\mathcal{C}}_{g,[n]},\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathfrak{u}}_{g,[n],\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
Applying the functor $H_{1}$ to the Lie algebra diagram above, we obtain the following comparison between the ordered and unordered cases.
Lemma 5.4.
For $g\geq 2$, $n\geq 1$, and $\ell\neq p$ an odd prime, there is a commutative diagram of $\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell}})$-representations with exact rows:
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}({\mathfrak{p}}_{g})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}({\mathfrak{u}}_{{\mathcal{C}}_{g,n},\bar{\xi}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}({\mathfrak{u}}_{g,n,\bar{\xi}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}({\mathfrak{p}}_{g})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}({\mathfrak{u}}_{{\mathcal{C}}_{g,[n]},\bar{\xi}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}({\mathfrak{u}}_{g,[n],\bar{\xi}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
Proof.
The exactness of the top row
$0\longrightarrow H_{1}(\mathfrak{p}_{g})\longrightarrow H_{1}(\mathfrak{u}_{\mathcal{C}_{g,n},\bar{\xi}})\longrightarrow H_{1}(\mathfrak{u}_{g,n,\bar{\xi}})\longrightarrow 0$
follows from the comparison with the characteristic 0 case, using Proposition 5.2 and [LW25, Cor. 5.16].
For the bottom row, the unordered case, right exactness of the functor $H_{1}$ implies exactness at the right; it remains only to show that the map $H_{1}(\mathfrak{p}_{g})\to H_{1}(\mathfrak{u}_{\mathcal{C}_{g,[n]},\bar{\xi}})$ is injective. To this end, we consider the natural weight filtration $W_{\bullet}\mathfrak{p}_{g}$ arising from the weighted completion of $\pi_{1}(\mathcal{C}_{g,n/\mathbb{F}_{p}})$ (see [WAT19]), a variant of relative completion developed by Hain and Matsumoto (see [HM03, HAI11]).
A key point is that $H_{1}(\mathfrak{p}_{g})$ is pure of weight 1-1. Consequently, the weight filtration on $\mathfrak{p}_{g}$ coincides with its lower central series. Since the Hodge weight filtration on $\mathfrak{p}^{\mathrm{an}}_{g}$ is also known to coincide with its lower central series, the comparison isomorphism $\mathfrak{p}^{\mathrm{an}}_{g}\cong\mathfrak{p}_{g}$, which is induced by the isomorphism $\tilde{\lambda}:\pi_{1}(C_{\bar{\eta}})^{\mathrm{pro}\,\ell}\xrightarrow{\sim}\pi_{1}(C_{\bar{\xi}})^{\mathrm{pro}\,\ell}$ from Proposition 4.2, is strictly compatible with weights. This implies that the induced isomorphism on derivation algebras is an isomorphism of filtered Lie algebras:
$(\mathrm{Der}(\mathfrak{p}_{g}),W_{\bullet})\cong(\mathrm{Der}(\mathfrak{p}^{\mathrm{an}}_{g}),W_{\bullet}).$
Consider the commutative diagram induced by $\tilde{\lambda}$:
${\mathfrak{p}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Der}(\mathfrak{p}_{g})}$${\mathfrak{p}^{\mathrm{an}}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathrm{Der}(\mathfrak{p}^{\mathrm{an}}_{g})\ignorespaces\ignorespaces\ignorespaces\ignorespaces,}\scriptstyle{\cong}$
which yields the induced diagram on the weight 1-1 graded pieces:
${H_{1}(\mathfrak{p}_{g})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Gr}^{W}_{-1}\mathrm{Der}(\mathfrak{p}_{g})}$${H_{1}(\mathfrak{p}^{\mathrm{an}}_{g})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathrm{Gr}^{W}_{-1}\mathrm{Der}(\mathfrak{p}^{\mathrm{an}}_{g})\ignorespaces\ignorespaces\ignorespaces\ignorespaces.}\scriptstyle{\cong}$
Since the functor $\mathrm{Gr}^{W}$ is exact on the category of MHS, the map $\mathrm{Gr}^{W}_{-1}\mathfrak{p}^{\mathrm{an}}_{g}=H_{1}(\mathfrak{p}^{\mathrm{an}}_{g})\to\mathrm{Gr}^{W}_{-1}\mathrm{Der}(\mathfrak{p}^{\mathrm{an}}_{g})$ is injective. Consequently, the top map $H_{1}(\mathfrak{p}_{g})\to\mathrm{Gr}^{W}_{-1}\mathrm{Der}(\mathfrak{p}_{g})$ is also injective.
In the case over $\bar{\mathbb{Q}}$, the Lie algebra $\mathfrak{u}_{\mathcal{C}_{g,n}}$ is negatively weighted, i.e., $\mathfrak{u}_{\mathcal{C}_{g,n}}=W_{-1}\mathfrak{u}_{\mathcal{C}_{g,n}}$. Therefore, the conjugation action of $\mathfrak{u}_{\mathcal{C}_{g,n}}$ on $\mathfrak{p}^{\mathrm{an}}_{g}$ lands in $W_{-1}\mathrm{Der}(\mathfrak{p}^{\mathrm{an}}_{g})$. The comparison isomorphism from Proposition 4.2 induces the commutative diagram:
${\mathfrak{p}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathfrak{u}_{\mathcal{C}_{g,n},\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{ad}}$${\mathrm{Der}(\mathfrak{p}_{g})}$${\mathfrak{p}^{\mathrm{an}}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathfrak{u}_{\mathcal{C}_{g,n}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{ad}}\scriptstyle{\cong}$${\mathrm{Der}(\mathfrak{p}^{\mathrm{an}}_{g})\ignorespaces\ignorespaces\ignorespaces\ignorespaces.}\scriptstyle{\cong}$
It follows that the image of $\mathfrak{u}_{\mathcal{C}_{g,n},\bar{\xi}}$ lies in $W_{-1}\mathrm{Der}(\mathfrak{p}_{g})$. Furthermore, since the map $\mathfrak{u}_{\mathcal{C}_{g,n},\bar{\xi}}\to\mathfrak{u}_{\mathcal{C}_{g,[n]},\bar{\xi}}$ is surjective and the adjoint action factors through this quotient, we obtain a sequence of maps:
$\mathfrak{p}_{g}\longrightarrow\mathfrak{u}_{\mathcal{C}_{g,n},\bar{\xi}}\longrightarrow\mathfrak{u}_{\mathcal{C}_{g,[n]},\bar{\xi}}\longrightarrow W_{-1}\mathrm{Der}(\mathfrak{p}_{g}).$
Since the target space $\mathrm{Gr}^{W}_{-1}\mathrm{Der}(\mathfrak{p}_{g})$ is abelian, the composition induces the sequence on homology:
$H_{1}(\mathfrak{p}_{g})\longrightarrow H_{1}(\mathfrak{u}_{\mathcal{C}_{g,n},\bar{\xi}})\longrightarrow H_{1}(\mathfrak{u}_{\mathcal{C}_{g,[n]},\bar{\xi}})\longrightarrow\mathrm{Gr}^{W}_{-1}\mathrm{Der}(\mathfrak{p}_{g}).$
Because the composite map from $H_{1}(\mathfrak{p}_{g})$ to $\mathrm{Gr}^{W}_{-1}\mathrm{Der}(\mathfrak{p}_{g})$ was shown to be injective, the map $H_{1}(\mathfrak{p}_{g})\to H_{1}(\mathfrak{u}_{\mathcal{C}_{g,[n]},\bar{\xi}})$ must also be injective. ∎
5.5.2. Key exact sequences for the family $\mathcal{M}_{g,n}\to\mathcal{M}_{g}$
Next, since ${\mathfrak{p}}^{\mathrm{an}}_{g,n}$ has trivial center [NTU95, p. 201] and ${\mathfrak{p}}^{\mathrm{an}}_{g,n}\cong{\mathfrak{p}}_{g,n}$, the same holds for ${\mathfrak{p}}_{g,n}$ and for $F^{\mathrm{un}}_{g,n}$.
[…]
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F^{\mathrm{un}}_{g,n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{G}}_{g,n,\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{G}}_{g,\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${F^{\mathrm{un}}_{g,[n]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{G}}_{g,[n],\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{G}}_{g,\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
[…]
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F^{\mathrm{un}}_{g,n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{U}}_{g,n,\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{U}}_{g,\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${F^{\mathrm{un}}_{g,[n]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{U}}_{g,[n],\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{U}}_{g,\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
Passing to Lie algebras, we obtain a commutative diagram of pronilpotent Lie algebras with exact rows
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathfrak{p}}_{g,n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathfrak{u}}_{g,n,\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathfrak{u}}_{g,\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${{\mathfrak{p}}_{g,[n]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathfrak{u}}_{g,[n],\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathfrak{u}}_{g,\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
Applying the functor $H_{1}$ to the preceding diagram of pronilpotent Lie algebras yields the following commutative diagram of $\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell}})$-representations.
Lemma 5.5.
For $g\geq 2$, $n\geq 1$, and $\ell\not=p$ an odd prime, there is a commutative diagram of $\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell}})$-representations with exact rows
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}({\mathfrak{p}}_{g,n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}({\mathfrak{u}}_{g,n,\bar{\xi}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}({\mathfrak{u}}_{g,\bar{\xi}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}({\mathfrak{p}}_{g,[n]})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}({\mathfrak{u}}_{g,[n],\bar{\xi}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}({\mathfrak{u}}_{g,\bar{\xi}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
Proof.
Consider the extension
$1\longrightarrow\mathcal{U}_{g,n,\bar{\xi}}\longrightarrow\mathcal{G}_{g,n,\bar{\xi}}\longrightarrow\mathrm{Sp}(H_{\mathbb{Q}_{\ell}})\longrightarrow 1.$
The group $\mathcal{G}_{g,n,\bar{\xi}}$ acts by conjugation on the ordered fiber $F^{\mathrm{un}}_{g,n}$ and the unipotent radical $\mathcal{U}_{g,n,\bar{\xi}}$. These actions induce $\mathcal{G}_{g,n,\bar{\xi}}$-equivariant maps on the homology of the corresponding Lie algebras:
$H_{1}(\mathfrak{p}_{g,n})\longrightarrow H_{1}(\mathfrak{u}_{g,n,\bar{\xi}})\quad\text{and}\quad H_{1}(\mathfrak{u}_{g,n,\bar{\xi}})\longrightarrow H_{1}(\mathfrak{u}_{g,\bar{\xi}}).$
Since the action of $\mathcal{G}_{g,n,\bar{\xi}}$ on these abelianizations factors through $\mathrm{Sp}(H_{\mathbb{Q}_{\ell}})$, these maps are $\mathrm{Sp}(H_{\mathbb{Q}_{\ell}})$-equivariant.
For the unordered case, similarly, the conjugation action of $\mathcal{G}_{g,[n],\bar{\xi}}$ on the unipotent radicals $\mathcal{U}_{g,[n],\bar{\xi}}$ and $\mathcal{U}_{g,\bar{\xi}}$ induces a map on their Lie algebra homology:
$H_{1}(\mathfrak{u}_{g,[n],\bar{\xi}})\longrightarrow H_{1}(\mathfrak{u}_{g,\bar{\xi}}).$
This map is $\mathcal{G}_{g,[n],\bar{\xi}}$-equivariant. Since the action of $\mathcal{G}_{g,[n],\bar{\xi}}$ on the abelianization of its unipotent radical factors through $\mathrm{Sp}(H_{\mathbb{Q}_{\ell}})$, this induced map is $\mathrm{Sp}(H_{\mathbb{Q}_{\ell}})$-equivariant. Next, recall from Proposition 3.3 that we have the isomorphism
$H_{1}(F^{\mathrm{un}}_{g,[n]})\cong\left(H_{1}(F^{\mathrm{un}}_{g,n})\right)_{S_{n}}.$
The conjugation action of $\mathcal{G}_{g,n,\bar{\xi}}$ on $H_{1}(F^{\mathrm{un}}_{g,n})\cong H_{\mathbb{Q}_{\ell}}^{\oplus n}$ commutes with the permutation of the marked points (acting diagonally on the summands), so it descends to a well-defined action on $H_{1}(\mathfrak{p}_{g,[n]})\cong H_{1}(F^{\mathrm{un}}_{g,[n]})$. Consequently, the natural morphism
$H_{1}(\mathfrak{p}_{g,[n]})\longrightarrow H_{1}(\mathfrak{u}_{g,[n],\bar{\xi}})$
is $\mathcal{G}_{g,n,\bar{\xi}}$-equivariant. Since the induced action on both the source (the coinvariants) and the target factors through the symplectic representation, this morphism is $\mathrm{Sp}(H_{\mathbb{Q}_{\ell}})$-equivariant.
For the exactness of the top row, applying the continuous relative completion functor to the exact sequence of groups in Proposition 4.4 yields the commutative diagram:
${H_{1}(\mathfrak{p}_{g,n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}(\mathfrak{u}_{g,n,\bar{\xi}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}(\mathfrak{u}_{g,\bar{\xi}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}(\mathfrak{p}^{\mathrm{an}}_{g,n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${H_{1}(\mathfrak{u}_{g,n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${H_{1}(\mathfrak{u}_{g})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${0.}$
Here, the vertical isomorphisms are induced by the comparison isomorphisms provided by Proposition 4.4 and Proposition 5.2. A sketch of the argument for the exactness of the bottom row at the left is as follows. The Lie algebras $\mathfrak{p}^{\mathrm{an}}_{g,n}$, $\mathfrak{u}_{g,n}$, and $\mathfrak{u}_{g}$ carry natural weight filtrations $W$ induced by their underlying Mixed Hodge Structures [HAI97, Sec. 3]. Since the first homology $H_{1}$ is pure of weight 1-1, these weight filtrations coincide with the lower central series [HAI97, Lem. 4.7]. The map $\mathfrak{p}^{\mathrm{an}}_{g,n}\to\mathfrak{u}_{g,n}$ is injective because $\mathfrak{p}^{\mathrm{an}}_{g,n}$ is center-free. Since the functor $\mathrm{Gr}^{W}_{\bullet}$ is exact on the category of MHS, this injectivity is preserved on the graded pieces. In particular, the induced map on the weight 1-1 component,
$\mathrm{Gr}^{W}_{-1}\mathfrak{p}^{\mathrm{an}}_{g,n}=H_{1}(\mathfrak{p}^{\mathrm{an}}_{g,n})\longrightarrow H_{1}(\mathfrak{u}_{g,n})=\mathrm{Gr}^{W}_{-1}\mathfrak{u}_{g,n},$
is injective.
Since the bottom row is exact (as established above) and the vertical maps are isomorphisms, the top row is exact. Therefore, the map $H_{1}(\mathfrak{p}_{g,n})\to H_{1}(\mathfrak{u}_{g,n,\bar{\xi}})$ is also injective.
Applying the functor $H_{1}$ to the exact sequence
$\mathfrak{p}_{g,[n]}\longrightarrow\mathfrak{u}_{g,[n],\bar{\xi}}\longrightarrow\mathfrak{u}_{g,\bar{\xi}}\longrightarrow 0,$
we obtain the exact sequence of $\mathrm{Sp}(H_{\mathbb{Q}_{\ell}})$-representations:
$H_{1}(\mathfrak{p}_{g,[n]})\xrightarrow{j}H_{1}(\mathfrak{u}_{g,[n],\bar{\xi}})\xrightarrow{\pi}H_{1}(\mathfrak{u}_{g,\bar{\xi}})\longrightarrow 0,$
where we denote the projection by $\pi$ and the natural $\mathrm{Sp}(H_{\mathbb{Q}_{\ell}})$-equivariant map by $j$. By exactness, we have $\ker\pi=\mathrm{im}\,j$.
Recall from Proposition 3.3 that $H_{1}(\mathfrak{p}_{g,[n]})\cong H_{\mathbb{Q}_{\ell}}$. Since the standard representation $H_{\mathbb{Q}_{\ell}}$ is irreducible, the map $j$ must be either the zero map or injective.
Suppose for the sake of contradiction that $j$ is the zero map. Then $\mathrm{im}\,j=0$, which implies $\ker\pi=0$. Consequently, $\pi$ would be an isomorphism. However, this implies that $\mathrm{Hom}_{\mathrm{Sp}}^{\mathrm{cts}}(H_{1}(\mathfrak{u}_{g,[n],\bar{\xi}}),H_{\mathbb{Q}_{\ell}})=0$ by Proposition 5.2 and [HAI11, Thm. 9.11]. This contradicts Lemma 5.3. Thus, $j$ must be an injection. This completes the proof. ∎
Remark 5.6.
Consider the exact sequence for the ordered case (the top row of Lemma 5.5):
$0\longrightarrow H_{1}(\mathfrak{p}_{g,n})\longrightarrow H_{1}(\mathfrak{u}_{g,n,\bar{\xi}})\longrightarrow H_{1}(\mathfrak{u}_{g,\bar{\xi}})\longrightarrow 0.$
This is a short exact sequence of $\mathrm{Sp}(H_{\mathbb{Q}_{\ell}})$-representations. Since the symplectic group $\mathrm{Sp}(H_{\mathbb{Q}_{\ell}})$ is reductive, its category of finite-dimensional representations is semisimple. Consequently, every short exact sequence splits, providing an $\mathrm{Sp}(H_{\mathbb{Q}_{\ell}})$-equivariant isomorphism:
$H_{1}(\mathfrak{u}_{g,n,\bar{\xi}})\cong H_{1}(\mathfrak{p}_{g,n})\oplus H_{1}(\mathfrak{u}_{g,\bar{\xi}})\cong H_{{\mathbb{Q}}_{\ell}}^{\oplus n}\oplus H_{1}({\mathfrak{u}}_{g,\bar{\xi}}).$
This splitting descends to the unordered case (the bottom row of Lemma 5.5). The vertical maps in the diagram correspond to taking coinvariants under the action of the symmetric group $S_{n}$. Since the functor of taking coinvariants commutes with direct sums, and $S_{n}$ acts trivially on the term $H_{1}(\mathfrak{u}_{g,\bar{\xi}})$, we obtain:
$H_{1}(\mathfrak{u}_{g,[n],\bar{\xi}})\cong\left(H_{1}(\mathfrak{p}_{g,n})\right)_{S_{n}}\oplus\left(H_{1}(\mathfrak{u}_{g,\bar{\xi}})\right)_{S_{n}}\cong H_{1}(\mathfrak{p}_{g,[n]})\oplus H_{1}(\mathfrak{u}_{g,\bar{\xi}})\cong H_{{\mathbb{Q}}_{\ell}}\oplus H_{1}(\mathfrak{u}_{g,\bar{\xi}}).$
This provides an $\mathrm{Sp}(H_{\mathbb{Q}_{\ell}})$-decomposition for the unordered relative completion that is compatible with the $S_{n}$-action.
5.6. Proof of the main results
Finally, we prove Theorem 1 and Corollary 2.
Proof of Theorem 1.
[…]
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F^{\mathrm{un}}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{G}}_{{\mathcal{C}}_{g,n},\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{G}}_{g,n,\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{t}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F^{\mathrm{un}}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{G}}_{{\mathcal{C}}_{g,[n]},\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{G}}_{g,[n],\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{s}$${1.}$
Passing to Lie algebras, this yields a corresponding commutative diagram of pronilpotent Lie algebras:
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathfrak{p}}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathfrak{u}}_{{\mathcal{C}}_{g,n},\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathfrak{u}}_{g,n,\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{dt}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathfrak{p}}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathfrak{u}}_{{\mathcal{C}}_{g,[n]},\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathfrak{u}}_{g,[n],\bar{\xi}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{ds}$${0.}$
Applying Lemma 5.4, Lemma 5.5, and Remark 5.6, we obtain the commutative diagram
[…]
where $dt^{\mathrm{ab}}$ and $ds^{\mathrm{ab}}$ denote the abelianizations of $dt$ and $ds$, respectively.
By the comparison isomorphisms established in Proposition 5.2, the section $t$ induces a section $t_{\bar{{\mathbb{Q}}}}$ of $\mathcal{G}_{\mathcal{C}_{g,n}}\to\mathcal{G}_{g,n}$, and consequently an $\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell}})$-equivariant linear section $dt^{\mathrm{ab}}_{\bar{{\mathbb{Q}}}}$ of $H_{1}(\mathfrak{u}_{\mathcal{C}_{g,n}})\to H_{1}(\mathfrak{u}_{g,n})$. According to [LW25, Prop. 6.4], for $g\geq 3$, there are exactly $n$ such sections arising from a section of $\mathcal{G}_{\mathcal{C}_{g,n}}\to\mathcal{G}_{g,n}$. More precisely, after identifying via the comparison isomorphism, the restriction of $dt^{\mathrm{ab}}$ to the summand $H_{{\mathbb{Q}}_{\ell}}^{\oplus n}$ maps into the component $H_{1}(\mathfrak{p}_{g})\cong H_{{\mathbb{Q}}_{\ell}}$ (corresponding to the fiber $C_{\bar{\xi}}$) by the formula
$dt^{\mathrm{ab}}:(x_{1},\ldots,x_{n})\longmapsto x_{i}$
for some fixed $i\in\{1,\ldots,n\}$.
The remainder of the argument follows the proof of [LW25, Thm. 1]. It is shown there that such a map $dt^{\mathrm{ab}}$ (projection to a single coordinate) is not invariant under the action of $S_{n}$ for $n\geq 2$. This contradicts the commutativity of diagram (16), which requires the section to descend to the $S_{n}$-coinvariants. Therefore, no such section $s$ can exist. ∎
Proof of Corollary 2.
Every continuous section of the projection $\pi_{1}(\mathcal{C}_{g,[n]/k})\to\pi_{1}(\mathcal{M}_{g,[n]/k})$ induces a section of the geometric fundamental groups $\pi_{1}(\mathcal{C}_{g,[n]/\bar{k}})\to\pi_{1}(\mathcal{M}_{g,[n]/\bar{k}})$. By the universal property of continuous relative completion, this in turn induces a section of the relative completions $\mathcal{G}_{\mathcal{C}_{g,[n]},\bar{\xi}}\to\mathcal{G}_{g,[n],\bar{\xi}}$, and consequently an $\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell}})$-equivariant linear section of
$H_{1}(\mathfrak{u}_{\mathcal{C}_{g,[n]},\bar{\xi}})\longrightarrow H_{1}(\mathfrak{u}_{g,[n],\bar{\xi}}).$
By Theorem 1, no such section exists. Thus, the original projection admits no section. ∎
6. Notation and Glossary
Throughout this paper, $k$ denotes a field of characteristic $p>0$, and $\bar{k}$ its algebraic closure. Let $R:=W(\bar{k})$ be the ring of Witt vectors of $\bar{k}$, and let $K$ denote the fraction field of $R$. We write $\bar{\xi}$ for a geometric closed point of $\operatorname{Spec}R$ and $\bar{\eta}$ for a geometric generic point of $\operatorname{Spec}R$.
Let $L$ be an algebraically closed field of characteristic $p\geq 0$. For a stack $\chi$ over $L$, we denote by $\bar{\zeta}$ a geometric point of $\chi$.
All fundamental groups are étale fundamental groups, unless stated otherwise, for example $\pi_{1}^{\mathrm{top}}$ or $\pi_{1}^{\mathrm{orb}}$. All pro-algebraic groups and Lie algebras are defined over $\mathbb{Q}$$_{\ell}$.
Moduli stacks and universal curves
  • ${\mathcal{M}}_{g,n/k}$: the moduli stack of smooth proper genus-$g$ curves with $n$ ordered distinct marked points over $k$.
  • ${\mathcal{M}}_{g,[n]/k}:=[{\mathcal{M}}_{g,n/k}/S_{n}]$: the moduli stack of smooth proper genus-$g$ curves with $n$ unordered marked points.
  • $\pi_{g,n/k}:{\mathcal{C}}_{g,n/k}\to{\mathcal{M}}_{g,n/k}$: the universal curve in the ordered case.
  • $\pi_{g,[n]/k}:{\mathcal{C}}_{g,[n]/k}\to{\mathcal{M}}_{g,[n]/k}$: the universal curve in the unordered case.
Configuration spaces and their fundamental groups
Let $C$ be a smooth proper connected curve of genus $g$ over $\bar{k}$.
  • $F(C,n):=C^{n}\setminus\Delta$: the ordered configuration space of $n$ distinct points on $C$.
  • $F(C,[n]):=F(C,n)/S_{n}$: the unordered configuration space.
  • $F_{g,n}:=\pi_{1}(F(C,n))$: the algebraic fundamental group of the ordered configuration space.
  • $F_{g,[n]}:=\pi_{1}(F(C,[n]))$: the algebraic fundamental group of the unordered configuration space.
$\ell$-adic unipotent completions
  • $F_{g,n}^{\mathrm{un}}$, $F_{g,[n]}^{\mathrm{un}}$: the $\ell$-adic unipotent completions of $F_{g,n}$ and $F_{g,[n]}$, respectively.
  • ${\mathfrak{p}}_{g,n}:=\operatorname{Lie}(F_{g,n}^{\mathrm{un}})$,  ${\mathfrak{p}}_{g,[n]}:=\operatorname{Lie}(F_{g,[n]}^{\mathrm{un}})$: the associated pronilpotent Lie algebras.
  • ${\mathfrak{p}}_{g}:={\mathfrak{p}}_{g,1}$: the Lie algebra associated with the unipotent completion of $\pi_{1}(C)$.
Continuous relative completions
Let
$\rho_{{\mathbb{Q}}_{\ell}}:\pi_{1}({\mathcal{M}}_{g,n/\bar{k}},\bar{\xi})\longrightarrow\operatorname{Sp}(H_{{\mathbb{Q}}_{\ell}})$
denote the $\ell$-adic monodromy representation, where
$H_{{\mathbb{Q}}_{\ell}}:=H^{1}_{\mathrm{\acute{e}t}}(C,{\mathbb{Q}}_{\ell}(1))$
[…]
The same construction applies verbatim over any algebraically closed base field
$L\in\{\bar{k},\ \bar{K},\ \bar{{\mathbb{Q}}}\},$
with respect to a chosen geometric point $\bar{\zeta}$ of the corresponding stack. Unless explicitly indicated otherwise, we use the same notation for the continuous relative completions over each of these bases, specifying the geometric point only when it is relevant (e.g. in specialization arguments).
  • ${\mathcal{G}}_{g,n,\bar{\zeta}}$, ${\mathcal{G}}_{g,[n],\bar{\zeta}}$: the continuous relative completions of $\pi_{1}({\mathcal{M}}_{g,n/L},\bar{\zeta})$ and $\pi_{1}({\mathcal{M}}_{g,[n]/L},\bar{\zeta})$ with respect to the induced $\ell$-adic monodromy representation $\rho_{{\mathbb{Q}}_{\ell}}$.
  • ${\mathcal{G}}_{{\mathcal{C}}_{g,n},\bar{\zeta}}$, ${\mathcal{G}}_{{\mathcal{C}}_{g,[n]},\bar{\zeta}}$: the continuous relative completions of the fundamental groups of the ordered and unordered universal curves over $L$, obtained by composing $\rho_{{\mathbb{Q}}_{\ell}}$ with the natural homomorphisms induced by the universal curve.
  • ${\mathcal{U}}_{(\cdot)}$: the unipotent radical of a continuous relative completion.
  • ${\mathfrak{u}}_{(\cdot)}:=\operatorname{Lie}({\mathcal{U}}_{(\cdot)})$: the associated pronilpotent Lie algebra.
Remark 6.1.
In the characteristic 0 setting of our prequel [LW25], continuous relative completions were denoted by ${\mathcal{G}}^{{(\ell)}}$ in order to distinguish them from relative completions over $\mathbb{Q}$}. In the present paper, all relative completions are continuous and defined over $\mathbb{Q}$$_{\ell}$, and we therefore omit the superscript $(\ell)$} to simplify notation.